### v1

# Rethinking Whole-Body CT Image Interpretation: An Abnormality-Centric Approach

Abstract
Automated interpretation of CT images—particularly localizing and describing abnormal findings across multi-plane and whole-body scans—remains a significant challenge in clinical radiology. This work aims to address this challenge through four key contributions: (i) On taxonomy, we collaborate with senior radiologists to propose a comprehensive hierarchical classification system, with 404 representative abnormal findings across all body regions; (ii) On data, we contribute a dataset containing over 14.5K CT images from multiple planes and all human body regions, and meticulously provide grounding annotations for over 19K abnormalities, each linked to the detailed description and cast into the taxonomy; (iii) On model development, we propose OminiAbnorm-CT, which can automatically ground and describe abnormal findings on multi-plane and whole-body CT images based on text queries, while also allowing flexible interaction through visual prompts; (iv) On benchmarks, we establish three representative evaluation tasks based on real clinical scenarios. Through extensive experiments, we show that OminiAbnorm-CT can significantly outperform existing methods on all the tasks and metrics.
1 Introduction
Computed Tomography (CT) imaging has become a cornerstone of modern medicine, with over 3 billion scans performed annually around the world. The interpretation of these scans—particularly the generation of radiology reports—plays a crucial role in guiding clinical decision-making. However, systematically identifying and characterizing abnormal findings in whole-body CT images remains a cognitively demanding and time-consuming task for radiologists rao2025multimodal .
In the recent literature, artificial intelligence has started to make significant progress in CT image interpretation, particularly along two directions: first, AI models have been developed for organ-level segmentation Totalsegmentator ; jaus2023towards , that enable precise anatomical labeling and organ-grounded report generation zhang2024radgenome ; chen2025large . Yet a critical gap remains: segmenting organs alone is insufficient for clinical utility. What ultimately matters to clinicians are abnormal findings—lesions and any anomalies that inform diagnosis and treatment; second, models for automatic report generation have been trained on the recently introduced datasets with paired CT scans and corresponding radiology reports, describing clinical findings and impressions chen2024bimcv ; hamamci2024foundation . However, these datasets are typically limited to specific anatomical regions—most commonly chest CT—and often lack explicit visual grounding, which reduces explainability and increases the risk of hallucinated content. To address the above-mentioned challenges, we aim to explore the potential to develop a system that automatically detects, localizes, and describes all abnormal findings across multi-plane, whole-body CT images. This paradigm shift moves beyond regional reporting and anatomical segmentation, aligning AI interpretation more closely with the practical demands of radiological diagnosis.
Figure 1: OminiAbnorm-CT-14K is the first large-scale dataset for abnormal findings grounding and description across multi-plane and whole-body CT images. It covers 349 representative abnormal findings from 82 anatomical structures and 40 major systems or organs.
One crucial barrier to developing such a system is the absence of a unified, clinically meaningful taxonomy of abnormalities, which presents a major challenge for defining task scope and establishing continuous benchmarks. In this paper, we collaborated with 7 radiologists from 3 centers (each with 10-16 years of experience) to construct a detailed hierarchical taxonomy covering 404 representative abnormal findings, that are organized across 40 major anatomical regions and 82 sub-regions.
Building on this foundation, we introduce OminiAbnorm-CT-14K-the first large-scale dataset designed for abnormality grounding and description on multi-plane whole-body CT imaging. It comprises 14.5K CT images from Radiopedia radiopaedia , covering axial, coronal, and sagittal planes and diverse anatomical regions. All images and the paired reports have been rigorously reviewed by experienced radiologists. We invite 4 radiologists (with at least 7 years’ experience) to manually annotate around 19K abnormal findings on the images in the format of either bounding box or segmentation masks. All regional annotations are linked to the corresponding report descriptions, and further cast into our proposed taxonomy. These annotations cover 349 out of the 404 representative abnormal findings (86%) defined in our taxonomy across whole-body CT images. This dataset addresses a critical gap in the field, by providing high-quality, clinically relevant annotations at scale—overcoming limitations in existing datasets caused by privacy constraints and the high cost of expert annotation. All our annotations will be released to the community, while the images are accessible for research purposes after applying and paying the required fee to Radiopaedia [^1] https://radiopaedia.org/licence#data-sets.
Leveraging OminiAbnorm-CT-14K, we develop OminiAbnorm-CT, a novel system for grounded CT image interpretation. By bridging a multi-modal language model with a specialized segmentation module in a bidirectional workflow, it enables dynamic invocation of the segmentation module during the text generation, and grounded abnormality interpretation based on the segmentation results. In contrast to prior approaches—which often lack grounding abilities, generalizability across planes and body regions, or flexibility in prompt handling—OminiAbnorm-CT enables automatic localization and description of abnormalities across multi-plane, whole-body CT scans. More importantly, OminiAbnorm-CT is designed to robustly interpret diverse textual references and visual prompts, enabling interactive and clinically adaptive use cases aligned with radiologists’ workflows.
For a comprehensive evaluation, we define three clinically relevant tasks: (i) grounded report generation, (ii) text-guided grounded report generation, and (iii) visual-prompted report generation. Experiment results show that OminiAbnorm-CT significantly outperforms existing baselines across all tasks, marking a substantial step toward explainable, abnormality-centric CT imaging interpretation.
In summary, our contributions to advance abnormality-centric grounded CT image interpretation can be summarized as: (i) We propose a comprehensive taxonomy for abnormalities on CT images, presented in Section 2; (ii) We build OminiAbnorm-CT-14K, the first dataset for abnormality grounding and description across multi-plane whole-body CT images, detailed in Section 3; (iii) We develop a novel system OminiAbnorm-CT, for automatically localizing and reporting abnormalities, illustrated in detail in Section 4; (iv) We set up a comprehensive benchmark with three clinically representative tasks, and demonstrate the superiority of OminiAbnorm-CT in Section 5.
2 Taxonomy for Abnormal Findings
Given the diversity of abnormal findings, their complex relationships and synonyms, we first construct a comprehensive taxonomy system, to categorize all the abnormalities into representative classes.
Our taxonomy is guided by four fundamental principles to ensure scientific rigor, systematic organization, and clinical utility: (i) comprehensive coverage. It should include all the clinically significant abnormalities, encompassing both lesions and the abnormal changes on anatomical structures; (ii) image-based definition. Each abnormal finding is precisely characterized by its observable features on the radiology scan, such as physical properties and morphological characteristics, that form the essential foundation for diagnosis; (iii) hierarchical organization. Findings are grouped by clinically important anatomical structures and major systems or organs, aligning with standard image interpretation protocols. While cross-regional structures (like skeletal systems) remain independent groups for conciseness; (iv) modular design. Each category represents a fundamental, independent finding that can be combined with others to express complex and multifaceted abnormalities.
The taxonomy is developed through the following three stages, involving 7 radiologists with 10 to 16 years of experience from three clinical centers, (i) we establish an anatomical hierarchy framework based on standard atlases and classic textbooks for human anatomy gray2008gray ; (ii) the radiologists then systematically catalog the abnormal findings under each anatomical structure, based on their clinical experience and several authoritative textbooks adam2020grainger ; haaga2016ct ; Mandell2013CoreRadiology on medical imaging and radiological pathology; lastly, (iii) the radiologists conduct rigorous cross-validation and discussion to establish precise consensus on the definitions of controversial or ambiguous abnormalities.
We finally identified 404 representative abnormalities across 82 anatomies throughout the human body, organized within 40 major organs and systems. The details can be found in Appendix Section A.11.
3 Dataset Curation
In this section, we present OminiAbnorm-CT-14K, which establishes a data foundation for abnormality grounding and interpretation across whole-body CT images. The dataset construction pipeline is illustrated in Figure 2. Here, we start with an overview of the limitations of existing datasets in Section 3.1, which motivate the dataset construction. We then describe the data source in Section 3.2, detail the annotation procedure in Section 3.3, outline the extension to instruction-style data in Section 3.4, and conclude with a summary in Section 3.5.
3.1 Limitation in Existing Datasets
We analyze several widely used public datasets relevant to our task (summarized in Appendix Table 6) and identify the following limitations: (i) limited annotation coverage. Most datasets provide grounding annotations for only a small subset of lesion types. Despite efforts to map existing labels to our proposed taxonomy, the majority of abnormal finding categories remain unannotated; (ii) restricted imaging scope. Existing datasets are primarily limited to axial-plane CT images or focus on specific anatomical regions, overlooking the growing clinical need for interpretation across multi-plane, whole-body CT scans; (iii) insufficient image-report alignment. Many datasets lack explicit associations between localized abnormalities in the images and their corresponding descriptions in radiology reports, hindering the development of grounded, explainable systems. These limitations make existing datasets inadequate for benchmarking or developing models capable of automatic detection, localization, and description of abnormal findings in multi-plane, whole-body CT imaging.
3.2 Data Source
Following prior works wu2023towards ; zheng2024large ; wu2024mrgen , we collect raw data from Radiopaedia radiopaedia , a publicly accessible platform of peer-reviewed clinical cases contributed by clinicians worldwide, with appropriate privacy safeguards. After application, we have obtained proper permission for all data used in this study. Shown in Figure 2(a), each case includes radiology scans across various planes and anatomical regions, along with patient background information and detailed clinical reports describing findings and diagnoses. In total, we collected 46,721 CT images paired with 14,920 clinical reports.
Figure 2: Data curation overview. (a) The CT image-report pairs are collected from an open-sourced and expert-checked website; (b) and (c) Radiologists provide grounding annotation on any abnormal findings, link to their text description in reports, and categorize into the taxonomy devised by a group of senior radiologists. The annotated data are further extended to instruction data with simulated visual prompts and text queries.
3.3 Annotation Pipeline
As illustrated in Figure 2(b) and (c), our goal is to produce grounding annotations for all abnormal findings in the collected CT images, i.e., linking them to corresponding textual descriptions in clinical reports, and categorizing them using our proposed taxonomy. To reduce annotation burden, we adopt a two pre-processing strategy: (i) image-side pre-processing. Rather than annotating entire scans, we focus on the key slice and its 8 adjacent slices, as identified and verified by radiologists to be the most informative and representative; (ii) text-side pre-processing. We employ GPT-4o gpt4o to extract findings from the reports and identify abnormalities using RaTEScore zhao2024ratescore .
Here, we recruited 4 clinicians with at least 7 years of radiology experience as annotators. To ensure high-quality annotations, we followed three core principles: (i) representative slices. Select the most representative slice(s) for annotation when abnormalities are present; (ii) complete annotation. Annotate all abnormal findings visible on selected slices; (iii) vision-language consistency. Only annotate abnormalities identifiable in provided slices; similarly, linked descriptions must reflect findings visible in these slices. For example, when a report mentions organ enlargement, the finding should not be annotated if enlargement cannot be confidently determined from the available slices.
Throughout the annotation process, we continuously refined OminiAbnorm-CT-14K based on annotator feedback: (i) taxonomy refinement. We remove or merge 4 less representative categories, add 22 previously omitted but clinically important ones, and revise 21 categories for improved clarity and expression; (ii) long-tail mitigation. To address data imbalance, we identify the underrepresented categories and use GPT-4o to filter scans referencing these rare findings, prioritizing them for annotation. As shown in Appendix Section A.6, it has significantly mitigated the long-tail distribution.
3.4 Extension to Instruction Data
As shown in Figure 2(b), we extend our annotated data to support three instruction tasks that simulate realistic clinical scenarios: (i) grounded report generation simulates a comprehensive examination. We format instructions to detect, ground, and describe all abnormal findings on CT images; (ii) text-guided grounded report generation involves responding to text queries about specific abnormalities of interest. For instance, for a patient with tuberculous empyema history, checking any abnormality related to it. We generate these queries using GPT-4o based on clinical reports, patient presentations, and medical history, simulating how radiologists approach images with prior knowledge; (iii) visual prompted report generation focuses on interpreting the marked abnormality. We create visual prompts from annotation masks using various formats, e.g., bounding box, ellipse, contour, and cropped region. This simulates a semi-automated workflow where the model elaborates the clinician-identified abnormality in details. More details are in Appendix Section A.10.
3.5 Summary
We prioritize axial CT images for annotation, while also incorporating coronal and sagittal images based on clinical recommendations. In total, we annotated 18,969 abnormalities on 9,990 axial scans, 2,738 coronal scans, and 1,803 sagittal scans. Representative examples are shown in Figure 1, and the distribution of annotated abnormalities is detailed in Appendix Section A.11. Annotation quality is verified by a senior radiologist with 12 years’ experience, with details in Appendix A.5.
4 OminiAbnorm-CT
This section presents the details of OminiAbnorm-CT, a novel system for grounded CT image interpretation . We start with the problem formulation in Section 4.1, then its architectural details in Section 4.2 and finally the training details in Section 4.3.
4.1 Problem Formulation
Given a CT image ($\mathcal{I}\in\mathbb{R}^{H\times W\times D}$), this paper aims to build a model $\Phi_{\theta}(\cdot)$, that enables generating the textual descriptions for abnormal regions, with intermediate groundings when necessary:
$\{\mathcal{R},\mathcal{S}\}=\Phi_{\theta}(\mathcal{T},\mathcal{I},\mathcal{V}),$ (1)
where $\mathcal{T}$ is the text instruction, $\mathcal{V}\in\mathbb{R}^{H\times W}$ (optional) is the visual prompt on a key slice from the users, for example, bounding boxes, ellipse, contour. $\mathcal{R}$ denotes the generated description for abnormalities, and $\mathcal{S}\in\mathbb{R}^{H\times W}$ refers to the (optional) grounding results on the key slice.
In contrast to the conventional approaches, $\Phi_{\theta}(\cdot)$ can interpret user-provided textual queries and ground arbitrary abnormal findings across multi-plane CT images spanning the entire human body, and generate corresponding descriptions. Additionally, it also supports visual prompts that allow the model to elaborate on user-highlighted abnormalities, enabling an interactive workflow where users can iteratively refine or correct the grounding results for more precise and context-aware descriptions.
4.2 Architecture
Figure 3: OminiAbnorm-CT. We bridge a multi-modal language model and a segmentation module, to allow grounding evidence acquisition during the generation of abnormality description, and further enhance its comprehension for flexible usage with text instruction and visual prompts.
The overall architecture of OminiAbnorm-CT is illustrated in Figure 3, which comprises a standard multi-modal language model and a segmentation module. Our core innovation lies in the integration of them, which enables localizing the abnormality and generating a textual description (findings), supporting flexible interaction through both visual prompts and textual instructions.
Multi-modal language model functions as the core to comprehend user instruction, invoke the segmentation module properly, and document the abnormal findings based on the grounding evidence. Specifically, in one workflow to generate grounded reports, given a CT image ($\mathcal{I}$) and text instruction ($\mathcal{T}$) as user input, it will first reason for the specific abnormal findings of interest for the user (or simply any abnormalities). For instance, when referring to patient history of tuberculous empyema, it should focus on related abnormalities including pleural thickening, calcifications, etc.
Then, to let it acquire grounding evidence, we expand the vocabulary with special token <SEG> as grounding request, which will invoke the segmentation module during generation. We extract its hidden state $h_{\text{seg}}$ from the last decoder layer as the prompt for the segmentation module, which encapsulates the targeted abnormal findings:
$\mathcal{S}=\Phi_{\text{seg}}(\mathcal{I},h_{\text{seg}}),\text{ }h_{\text{seg}}\in\mathbb{R}^{d}$ (2)
Given the segmentation results ($S$), we convert it to visual prompt ($\mathcal{V}$) in box format, i.e., directly overlay it on the original image to create a new input image ($\mathcal{I}^{\prime}$), and the multi-modal language model could seamlessly continue generating the abnormality description ($\mathcal{R}$) based on it:
$\mathcal{R}=\Phi_{\text{MLLM}}(\mathcal{T},\mathcal{I}^{\prime}),\text{ }\mathcal{I}^{\prime}=\mathcal{I}\oplus\mathcal{V}$ (3)
where $\oplus$ represents the superimposing operation. This leverages the inherent ability of multi-modal language model to perceive visual markers cai2024vip , guiding its attention to specific regions of interest.
Note that, the above design naturally supports another workflow when users have manually delineated specific abnormalities, or intend to refine the grounding results. In such case, $\mathcal{V}$ can also be user-input visual prompts on the image, and the multi-modal language model elaborates the user-highlighted abnormality following Equation 3.
Segmentation module is invoked for abnormality grounding based on the CT image ($\mathcal{I}$) and prompt ($h_{\text{seg}}$), summarized as Equation 2. Specifically, it first adopts an encoder-decoder backbone to derive image features:
$(v,u)=\Phi_{\text{seg}}(\mathcal{I}),\text{ }v\in\mathbb{R}^{H^{\prime}\times W^{\prime}\times d^{\prime}},\text{ }u\in\mathbb{R}^{H\times W\times d},$ (4)
where $v$ denotes concatenation of multi-scale image embeddings along the channel dimension, which are down-sampled from each encoder layer to a unified resolution, and $u$ is the pixel-level dense feature from the last decoder layer. To bridge the gap between latent spaces, we further align the segmentation prompt ($h_{\text{seg}}$) with the image embeddings ($v$):
$q=\Phi_{\text{crossattn}}(v,f(h_{\text{seg}})),\text{ }q\in\mathbb{R}^{d}$ (5)
where $f$ is a projection layer for dimension consistency, and $\Phi_{\text{crossattn}}$ is a cross-attention module treating $v$ as key and value, $f(h_{\text{seg}})$ as query. The segmentation prediction is then derived by performing dot product between the adapted feature $q$ and $u$:
$\mathcal{S}=q\cdot u,\text{ }\mathcal{S}\in\mathbb{R}^{H\times W},$ (6)
4.3 Training
Training object. We train the multi-modal language model and the segmentation module jointly with a text generation loss and a segmentation loss:
$\mathcal{L}_{\textbf{txt}}=\text{CE}(\hat{\mathcal{R}},\mathcal{R}),\text{\hskip 6.0pt}\mathcal{L}_{\textbf{seg}}=\text{BCE}(\hat{\mathcal{S}},\mathcal{S})+\text{DICE}(\hat{\mathcal{S}},\mathcal{S})$ (7)
Training data. Our training data primarily originates from OminiAbnorm-CT-14K, and is formulated into three instruction data formats as illustrated in Section 3.4. Additionally, we incorporate two types of data as supplements: (i) lesion segmentation data. This type of data can be formulated into grounding tasks and utilized to enhance the grounding ability. A detailed list of these datasets can be found in Appendix Table 6; (ii) medical VQA data. This type of data is involved to maintain our model’s generalization capabilities with more diverse question types. Specifically, we take CT images and corresponding question-answering data from PubMedVision chen2024huatuogpt .
Implementation details. In practice, we adopt QWen2.5-VL-7B yang2024qwen2 as the multi-modal language model, and a 6-layer U-Net as the segmentation module. The cross-attention module consists of 6 layers. As shown in Appendix Section A.8, we find that expanding the input context range (i.e., using more slices) does not enhance the report generation performance of the multi-modal language model. Thus, we only feed the central slice to the model. However, for the segmentation module, we include up to 8 adjacent slices as additional input. To save computational cost, we only optimize the inserted LoRA layers hu2022lora in the multi-modal language model while leaving the whole segmentation module trainable. More implementation details are in Appendix Section A.3.
5 Experiments
In this section, we first present experimental settings in Section 5.1, including task formulation, metrics, baselines, and test datasets. We then provide detailed quantitative experimental results in Section 5.2, and ablation study in Section 5.3. In addition, the qualitative experimental results are in the Appendix Section A.1 and more ablation studies are in the Appendix A.8 and A.9.
Table 1: Quantitative results on visual prompted report generation. Results are averaged within each category and then across all categories, with the best bolded and the second best underlined. Models optimized with medical data are marked with ✓.
Model B-1 B-2 B-3 RTSc BTSc MTR R-1 R-L RG
Axial (n=2193)
GPT-4o 15.80 6.15 1.59 39.04 84.58 19.31 18.62 14.41 6.51
QWen2.5-VL-7B 11.57 4.20 0.76 38.35 84.10 16.68 14.26 11.09 3.85
ViP-LLaVA 15.21 4.49 0.69 33.07 85.15 13.65 17.41 14.68 3.35
MedFlamingo ✓ 11.26 1.71 0.00 17.35 80.34 9.43 11.69 10.55 0.50
MedDr ✓ 9.54 2.46 0.41 32.24 78.80 9.13 12.23 10.34 3.74
LLaVA-Med ✓ 15.55 4.23 0.55 39.08 84.68 15.58 17.67 13.55 6.06
BiomedGPT ✓ 10.80 4.16 0.71 38.16 83.19 15.77 13.40 10.81 4.65
OminiAbnorm-CT ✓ 18.03 6.65 1.80 42.81 86.35 19.40 21.66 17.61 9.98
Coronal (n=750)
GPT-4o 9.34 3.15 0.96 27.13 58.83 9.27 12.25 10.01 4.31
QWen2.5-VL-7B 13.62 3.54 0.66 37.26 84.19 13.49 14.97 11.47 3.37
ViP-LLaVA 13.85 3.81 0.53 35.21 84.60 12.52 17.07 13.96 3.80
MedFlamingo ✓ 10.98 1.66 0.00 17.73 79.98 9.01 11.60 10.34 0.53
MedDr ✓ 8.07 2.07 0.29 32.77 76.07 8.09 11.55 9.63 4.08
LLaVA-Med ✓ 15.65 4.18 0.68 40.99 84.38 14.18 18.15 13.52 6.54
BiomedGPT ✓ 12.60 4.79 0.70 38.47 82.89 15.45 14.78 11.52 4.45
OminiAbnorm-CT ✓ 18.38 7.19 2.39 42.88 86.00 19.00 21.30 16.87 9.44
Sagittal (n=591)
GPT-4o 10.29 3.40 0.84 32.90 72.76 10.07 13.86 11.63 4.04
QWen2.5-VL-7B 13.28 3.54 0.63 35.67 83.98 13.01 14.75 11.34 2.91
ViP-LLaVA 12.84 3.83 0.38 34.51 84.44 11.88 16.29 13.44 2.86
MedFlamingo ✓ 10.57 1.53 0.00 17.36 78.87 8.66 11.40 10.14 0.36
MedDr ✓ 6.98 1.77 0.28 30.77 75.35 7.12 10.32 8.54 2.42
LLaVA-Med ✓ 15.10 4.68 0.66 39.19 84.10 14.18 17.90 13.57 5.11
BiomedGPT ✓ 12.59 4.81 0.73 36.62 82.60 15.19 14.62 11.35 3.37
OminiAbnorm-CT ✓ 17.29 6.48 1.82 41.77 85.62 17.45 20.43 15.87 7.51
5.1 Experiment Settings
Based on OminiAbnorm-CT-14K, we build a rigorous and comprehensive benchmark to evaluate both grounding and description generation abilities for abnormalities on whole-body CT images.
Task formulation. As detailed in Section 3.4, we evaluate methods on three tasks of increasing difficulty: visual prompted report generation, grounded report generation, and text-guided grounded report generation. For text-guided grounded report generation, we generate queries for both present and absent abnormalities at a 3:1 ratio to simulate realistic diagnostic scenarios.
Metrics. For grounding results, we calculate Dice Similarity Coefficient (DSC); for generated descriptions, we calculate BLEU (B) bleu , RaTEScore (RTSc) zhao2024ratescore , BERTScore (BTSc) zhangbertscore , METEOR (MTR) banerjee2005meteor , ROUGE (R) rouge and RadGraph (RG) delbrouck2024radgraph . In the text-guided grounded report generation, we prompt models to output a specific response when the queried abnormalities are confirmed absent. Specificity (SP) is then calculated as the proportion of correctly identified absent abnormalities out of all truly absent abnormalities in the test set. Meanwhile, when models miss the queried abnormalities, both DSC and all description metrics are assigned as 0.
Baseline. For visual prompted report generation, we compare with multi-modal LLMs in both the medical domain (MedDr meddr , BiomedGPT zhang2024generalist , LLaVA-Med li2023llava , and Med-Flamingo moor2023med ) and the general domain (GPT-4o gpt4o , Qwen2.5-VL yang2024qwen2 , and ViP-LLaVA cai2024vip ). For (text-guided) grounded report generation, as no such solution exists in the literature, we prompt LLaVA-Med to describe abnormal findings based on segmentation results from auxiliary grounding models. Since the images in OminiAbnorm-CT-14K lack the original DICOM metadata, we only consider 2D segmentation models as baselines, including BiomedParse zhao2024foundation , LiSA lai2024lisa , and MedULS de2024uls23 . More details are in Appendix Section A.4.
Data. We independently split the train and test data for axial, coronal, and sagittal images. To guarantee diversity, we allocate at least 5 samples per category to the axial test set (2 for coronal and sagittal), or all samples if fewer exist, with the remaining data following a 3:1 train-test ratio. To avoid data leakage, images from the same patient are restricted to the same set.
5.2 Experiment Results
Table 2: Quantitative results on grounded report generation. Results are averaged within each category and across all categories, with the best results bolded and the second best underlined.
Model DSC B-1 B-2 B-3 RTSc BTSc MTR R-1 R-L RG
Axial (n=2193)
LiSA + LLaVA-Med 20.96 12.94 6.65 0.73 31.38 82.07 19.20 14.62 12.44 2.79
BiomedParse + LLaVA-Med 15.78 13.61 6.98 0.84 31.50 82.36 19.84 15.28 13.02 2.74
MedULS + LLaVA-Med 15.24 12.70 6.43 0.74 31.05 84.27 18.74 14.67 12.42 2.31
OminiAbnorm-CT 36.04 18.50 11.75 7.90 33.24 86.10 21.83 21.65 18.75 3.78
Coronal (n=750)
LiSA + LLaVA-Med 17.09 11.48 5.85 0.64 33.28 82.27 16.23 14.36 12.12 2.74
BiomedParse + LLaVA-Med 14.55 11.71 5.95 0.68 32.89 81.80 16.88 14.75 12.57 3.35
MedULS + LLaVA-Med 12.77 11.64 5.88 0.64 32.25 84.11 16.45 14.71 12.34 2.96
OminiAbnorm-CT 31.65 19.13 11.86 7.81 35.47 85.58 20.66 22.14 18.26 4.55
Sagittal (n=591)
LiSA + LLaVA-Med 14.72 11.10 5.60 0.58 31.51 81.12 16.18 14.07 11.96 2.18
BiomedParse + LLaVA-Med 14.45 10.89 5.46 0.52 32.10 79.37 15.68 14.06 11.95 2.13
MedULS + LLaVA-Med 12.95 10.95 5.32 0.50 30.94 83.53 15.39 14.12 11.83 1.87
OminiAbnorm-CT 34.38 17.10 10.63 7.10 33.13 85.25 19.05 20.92 17.67 3.70
Visual Prompted Report Generation. As depicted in Table 1, OminiAbnorm-CT consistently outperforms all baselines across all planes and metrics. Specifically, compared to the strongest baseline LLaVA-Med, OminiAbnorm-CT achieves average improvements of 2.5 in BLEU-1, 2.7 in RaTEScore, and 3.1 in RadGraph across all three planes. These results demonstrate OminiAbnorm-CT’s effectiveness in semi-automated workflows, where radiologists identify abnormalities and the model provides accurate and detailed interpretations.
Grounded Report Generation. As shown in Table 2, OminiAbnorm-CT demonstrates superior grounding capabilities, outperforming the best baselines by 15.08, 14.56, and 19.66 in DSC across axial, coronal, and sagittal planes respectively. For report generation, OminiAbnorm-CT achieves the best performance across all metrics, with notable improvements in both RadGraph and RaTEScore across all anatomical planes, indicating that the generated reports can better align with clinical standards. These results demonstrate OminiAbnorm-CT’s potential for fully automated abnormality localization and report generation in clinical settings.
Table 3: Quantitative results on text-guided grounded report generation. Results are averaged within each category and then across all categories, with the best bolded and the second best underlined.
Model DSC SP B-1 B-2 B-3 RTSc BTSc MTR R-1 R-L RG
Axial (n=2193)
MedULS + LLaVA-Med 14.53 6.10 9.54 2.18 0.19 27.35 82.98 11.45 11.48 9.20 1.61
LiSA + LLaVA-Med 17.48 21.35 10.57 2.41 0.28 29.18 83.47 11.59 12.34 9.80 1.61
BiomedParse + LLaVA-Med 16.57 19.87 10.66 2.38 0.26 29.66 83.39 11.74 12.62 9.86 2.16
OminiAbnorm-CT 32.40 58.43 11.94 3.33 0.74 34.80 83.85 12.94 14.65 11.94 4.32
Coronal (n=750)
MedULS + LLaVA-Med 12.26 5.37 10.73 2.60 0.26 29.34 82.89 11.40 12.79 9.82 1.61
LiSA + LLaVA-Med 12.04 13.76 11.58 2.71 0.22 30.76 83.21 11.36 13.40 10.23 1.83
BiomedParse + LLaVA-Med 13.60 17.79 11.21 2.64 0.25 30.24 83.33 11.47 13.50 10.31 2.34
OminiAbnorm-CT 27.74 45.16 12.51 4.10 1.22 36.56 82.98 13.61 15.83 12.36 5.72
Sagittal (n=591)
MedULS + LLaVA-Med 11.89 7.66 10.25 2.69 0.29 27.40 82.49 11.24 12.17 9.53 1.23
LiSA + LLaVA-Med 11.23 16.67 11.14 2.72 0.30 29.79 82.88 11.15 13.36 10.07 1.42
BiomedParse + LLaVA-Med 15.35 17.57 11.07 2.72 0.39 31.12 83.01 11.66 13.85 10.40 1.78
OminiAbnorm-CT 29.89 45.41 12.80 4.51 1.42 36.04 83.97 14.02 16.20 12.79 4.85
Text-guided Grounded Report Generation. As shown in Table 3, OminiAbnorm-CT demonstrates superior grounding capabilities, with DSC scores improving by 14.24, 14.14, and 14.54 in axial, coronal, and sagittal planes respectively compared to the best baselines. For report generation, OminiAbnorm-CT excels across both lexical overlap metrics (average improvements of 1.3 on BLEU-1 and 2.2 on ROUGE-1 across all planes) and clinical relevance metrics (average gains of 4.0 on RaTEScore and 2.9 on RadGraph across all planes), confirming its superior ability to generate reports that align with clinical assessments. Additionally, OminiAbnorm-CT achieves significantly higher Specificity, demonstrating its robustness to queries about non-existing abnormalities.
5.3 Ablation Study
The integration of the segmentation module allows OminiAbnorm-CT to acquire grounding evidence for abnormal findings and generate descriptions accordingly. To validate its necessity, we conduct an ablation study comparing the full model against a variant where the segmentation module is removed (denoted as OminiAbnorm-CT w/o Seg). This modified version directly generates findings descriptions without anatomical localization guidance.
In the text-guided grounded report generation task, as shown in Table 4, OminiAbnorm-CT without the segmentation module demonstrates notable performance degradation, with inferior results on 23 of 27 metrics across three planes. Furthermore, detailed analysis reveals that without the segmentation module, the model struggles to accurately detect the queried abnormalities in CT images, exhibiting a pronounced negative bias that leads to high specificity scores.
Similarly, in the grounded report generation task, as depicted in Table 5, removing the segmentation module leads to consistent performance degradation on all the metrics, further confirming that segmentation-based evidence is crucial for generating high-quality medical reports.
Table 4: Ablation study of the segmentation module for text-guided grounded report generation task. Results are averaged within each category and then across all categories, with the best bolded
Model SP B-1 B-2 B-3 RTSc BTSc MTR R-1 R-L RG
Axial (n=2193)
OminiAbnorm-CT w/o Seg 95.39 4.50 2.33 1.21 20.53 31.06 7.04 8.50 7.02 7.25
OminiAbnorm-CT 58.43 11.94 3.33 0.74 34.80 83.85 12.94 14.65 11.94 4.32
Coronal (n=750)
OminiAbnorm-CT w/o Seg 95.71 4.50 2.47 1.27 19.07 28.34 6.27 7.96 6.43 6.17
OminiAbnorm-CT 45.16 12.51 4.10 1.22 36.56 82.98 13.61 15.83 12.36 5.72
Sagittal (n=591)
OminiAbnorm-CT w/o Seg 92.08 4.17 2.16 1.12 16.62 25.43 5.84 7.06 5.53 5.67
OminiAbnorm-CT 45.41 12.80 4.51 1.42 36.04 83.97 14.02 16.20 12.79 4.85
Table 5: Ablation study of the segmentation module in grounded report generation task. Results are averaged within each category and then across all categories, with the best bolded
Model B-1 B-2 B-3 RTSc BTSc MTR R-1 R-L RG
Axial (n=2193)
OminiAbnorm-CT w/o Seg 7.73 4.74 3.16 27.87 73.14 17.65 15.18 13.58 3.42
OminiAbnorm-CT 19.00 12.26 8.40 33.85 86.24 22.45 22.27 19.43 4.60
Coronal (n=750)
OminiAbnorm-CT w/o Seg 8.12 4.81 3.12 31.03 75.55 15.64 15.27 13.23 3.84
OminiAbnorm-CT 19.28 12.03 7.96 35.40 85.56 20.81 22.35 18.47 4.65
Sagittal (n=591)
OminiAbnorm-CT w/o Seg 8.79 5.21 3.33 31.21 77.41 16.18 15.86 13.87 3.91
OminiAbnorm-CT 16.69 10.41 6.98 33.45 84.94 18.77 20.67 17.61 3.83
6 Related Work
Medical image segmentation aims to delineate clinically meaningful regions on medical images. For radiology images, most datasets focus on annotations for important organs and anatomical structures Totalsegmentator ; abdomenatlas ; FLARE ; AMOS22 or lesions heller2021kidney ; MSD ; BraTS2023GLI ; de2024uls23 ; yan2018deep . Specialist segmentation models nnUNET ; UNet ; zhou2023nnformer ; zhou2019unet++ have achieved remarkable success in the past decade, while recently, building foundational generalist models MedSAM ; zhao2024foundation ; zhao2023one ; du2024segvol has received increased attention. However, few works contribute to the grounding of abnormal image findings, which encompass a broader spectrum including abnormal organs, potential lesions, and more generally, any anomalies meaningful for clinical decision-making.
Medical image report generation aims to faithfully interpret the medical images with formal text, which is indispensable for numerous clinic procedures rao2025multimodal . Existing datasets with image-report pairs are limited to specific modalities and regions, such as chest X-ray johnson2019mimic , brain MRI lei2024autorg and chest CT hamamci2024foundation ; chen2024bimcv , overlooking the clinical demands for multi-plane whole-body CT image report generation. Recent progress zhang2024radgenome ; chen2025large ; lei2024autorg ; chen2024vision on grounded report generation are limited to structured organ-level segmentation and description, failing to localize and interprete fine-grained abnormalities.
Generative visual-language foundation models for medicine are built on large-scale multi-modal medical data, and have demonstrated exceptional performance and generalization capabilities across diverse medical tasks wu2023towards ; moor2023med ; zhang2024generalist ; li2023llava ; meddr . Despite their potential application to multi-plane, whole-body CT images, however, most existing approaches fail to simultaneously provide grounded evidence during generation, leading to poor explainability and hallucinations that are more difficult to detect.
7 Conclusion
This paper advances the automatic grounded interpretation of CT imaging from an abnormality-centric view. To support this, we develop a hierarchical taxonomy of 404 representative abnormal findings on CT images. We contribute OminiAbnorm-CT-14K, a meticulously annotated dataset with detailed grounding and description annotation for around 19K abnormalities, from 14.5K multi-plane CT images across the entire human body. Built on this dataset, OminiAbnorm-CT enables automatic grounding and description of abnormalities, while supporting flexible clinical usage with text queries or visual prompts. Comprehensive evaluations across three clinically representative tasks demonstrate OminiAbnorm-CTs superior performance. Our work pioneers abnormality-centric CT interpretation, enhancing diagnostic transparency through accurate grounding and detailed characterization of findings, with significant potential to transform radiological practice.
Acknowledgments and Disclosure of Funding
Weidi would like to acknowledge the funding from Scientific Research Innovation Capability Support Project for Young Faculty (ZY-GXQNJSKYCXNLZCXM-I22).
Appendix A Appendix
A.1 Qualitative Experiment Results
Fig. 4 presents several examples from the grounded report generation task, comparing the segmentation and report generation results between OminiAbnorm-CT and BiomedParse+LLaVA-Med. The results clearly demonstrate that BiomedParse fails to detect any abnormalities in the latter two cases, consequently leading LLaVA-Med to generate reports that are hardly relevant to the images. In contrast, OminiAbnorm-CT successfully localizes the abnormal regions across all examples and produces comparatively more accurate reports for the abnormalities.
Figure 4: Qualitative comparison on the grounded report generation task.
In Fig. 5, we demonstrate several examples from the text-guided grounded report generation task, and compare the segmentation and report generation results of OminiAbnorm-CT and BiomedParse+LLaVA-Med. As illustrated in the first two examples, BiomedParse fails to accurately localize the queried abnormalities. While in the latter two examples, LLaVA-Med is unable to correctly interpret the segmentation results when generating reports. In contrast, OminiAbnorm-CT can successfully localize the queried abnormalities across all examples and generates more precise and clinically relevant reports.
Figure 5: Qualitative comparison on the text-guided grounded report generation task.
In Fig. 6, we demonstrate several representative examples in the visual prompted report generation task, comparing the abnormality descriptions generated by OminiAbnorm-CT, LLaVA-Med, and QWen2.5-VL-7B. It shows that LLaVA-Med and QWen2.5-VL-7B exhibit some fundamental errors in their results. For instance, in the first case, LLaVA-Med misidentifies the kidney as the gallbladder; While in the last case, QWen2.5-VL-7B fails to distinguish between the liver and gallbladder. By contrast, OminiAbnorm-CT is able to precisely capture and identify the abnormal findings highlighted by the visual prompts and generate high quality reports.
Figure 6: Qualitative comparison on the visual prompted report generation task.
A.2 Limitations
As a pioneering attempt toward automatic grounding and interpretation of abnormalities on multi-plane and whole-body CT images, our work certainly has several limitations: (i) Regarding data annotation, despite providing a large-scale dataset, some rare abnormalities (55 out of 404 in the taxonomy) remain uncovered, as shown in Section A.11. Additionally, many abnormalities have very few annotations, which significantly restricts the model’s capability range. Expanding annotations for these rare categories is treated as our future work. Meanwhile, the CT images in OminiAbnorm-CT-14K may also exhibit demographic and regional bias. (ii) Due to resource constraints, we currently provide grounding annotation for a few representative slices on each CT image, rather than full volumetric segmentation. In the future, we could potentially address this limitation by using interactive segmentation models such as nnInteractive [27]. (iii) Limited by the 3D image perception capability of multi-modal language models, which is also a common challenge in current generative visual-language foundation models for medicine [53, 44, 58, 33, 20], we only considered single slice as input when generating reports. Enhancing models’ ability to reason across adjacent slices and integrate volumetric context represents a critical direction for our future research.
A.3 More Implementation Details of OminiAbnorm-CT
Training Hyperparameters. We adopt 8 NVIDIA A100-80G GPUs for training, with 1 sample per device and gradient accumulation set to 4. We train OminiAbnorm-CT for 250K iterations in total, and use 5% steps for warming up. We take AdamW [36] as optimizer with a learning rate of 2e-4. When optimizing the multi-modal language model and the segmentation module jointly, we assign equal weights to the text generation loss and segmentation loss, as defined in Equation 7. We set the rank of LORA to 32.
Training Data. We pad and rescale all images to 512$\times$512. For training data sampling from four tasks: visual prompted report generation, grounded report generation, text-guided grounded report generation, and general VQA, we control the ratio to be 1:1:1:1. Additionally, in the abnormality grounding task, we maintain a ratio of 6:4 between OminiAbnorm-CT-14K and public lesion segmentation data. For each abnormality description, we use the following prompt to have GPT rewrite three different versions, then randomly select one as the ground truth:
Your task is to produce THREE different rewrites of medical reports while preserving all original medical content and diagnostic information, changing only the expression style and sentence structure. Rules: 1. Strictly maintain all medical information, diagnostic results, numerical values, and key findings. 2. Do not add any new medical content or conclusions. 3. Do not remove any medical information from the original report. 4. Only change the wording, sentence structure, word order, and vocabulary choices. 5. Maintain professional medical language and terminology. 6. Preserve the overall structure of the original report (such as section divisions). 7. Ensure all rewritten reports remain medically rigorous and accurate. 8. Create THREE distinct rewrites with different phrasing and structure. Input: Original medical report written by a doctor Output: THREE rewritten versions with identical content but different expression Please rewrite the following medical report in three different ways: \$Findings Description\$ Your response must follow this exact format: \$Output Template\$
A.4 Implementation Details of Baselines
In visual prompted generation task, we prompt each abnormality with all types of visual prompts, and take the maximum score. All the methods are prompted with the following template:
You are a helpful medical assistant. Describe the abnormal findings indicated by the \$Visual Prompt\$. Please use precise medical terminology, maintain the concise reporting style used in formal radiology reports and provide only the specific radiological findings. Do not list general possibilities, explanations, or recommendations.
In grounded report generation task, we integrate LLaVA-Med with 2D segmentation models as baselines. Specifically, we re-implement the MedULS [13] with a 2D nnU-Net [26] and the public lesion segmentation datasets in Table 6, covering all the datasets in the official ULS-23 challenge and 10 additional ones. To our knowledge, this represents the segmentation model with the broadest capability range (in terms of abnormality variety) that can be constructed from currently available public datasets. For BiomedParse [61] and LiSA [31], we prompt them with ‘Abnormal findings on the CT image’ to derive segmentation results for all the abnormalities on the input CT image. Then, we combine each segmentation model with LLaVA-Med by converting the segmentation results into bounding boxes overlaid on the CT image, and prompt LLaVA-Med to generate the report based on these visual cues:
You are a helpful medical assistant. The abnormal findings are highlighted in red boxes on this CT image, if present. Please describe each abnormal finding indicated by the red boxes using the format ’Finding 1: [description]’, ’Finding 2: [description]’, etc. Use precise medical terminology, maintain the concise reporting style used in formal radiology reports and provide only the specific radiological findings. Do not list general possibilities, explanations, or recommendations. Respond with ’I don’t see any abnormalities on the image.’ if no abnormalities are present.
In text-guided grounded report generation task, we use the simulated text queries as prompts for BiomedParse and LiSA to derive segmentation results for the queried abnormalities, which are detailed in Section 3.4. For MedULS, since it doesn’t support text-prompted segmentation, we simply use its unconditioned segmentation results. Similarly, we convert their segmentation results into bounding boxes overlaid on the CT image, and prompt LLaVA-Med to generate the report based on them:
You are a helpful medical assistant. The abnormal findings are highlighted in red boxes on this CT image, if present. Please describe the abnormal findings indicated by the red boxes. Use precise medical terminology, maintain the concise reporting style used in formal radiology reports and provide only the specific radiological findings. Do not list general possibilities, explanations, or recommendations. Respond with ’I don’t see any relevant abnormalities on the image.’ if no abnormalities are present.
A.5 Quality Verification on OminiAbnorm-CT-14K
We conducted rigorous and comprehensive quality verification on the annotations in OminiAbnorm-CT-14K. For each annotator, we randomly sampled 100 annotated images on the axial plane, 50 on the coronal, and 50 on the sagittal plane. Then a senior radiologist with 12 years of experience assesses the annotation quality on four key metrics: (i) Detection rate measures the percentage of abnormalities successfully identified and annotated by the annotator, without any omission; (ii) Grounding precision evaluates the percentage of grounding annotations that properly encompass the primary regions of the abnormality, while minimizing false positive areas; (iii) Report concordance quantifies the percentage of description annotations that faithfully reflect the linked abnormal findings on the image; (iv) Classification accuracy measures the percentage of abnormalities that are correctly categorized.
Our verification results demonstrated that: (i) 3 out of 4 annotators achieved perfect scores across all four metrics (100%). (ii) Only one annotator had 10 abnormality annotations where the descriptions were not sufficiently accurate and the category labels were also incorrect, resulting in Report concordance and Classification accuracy of 95%, while maintaining 100% Detection rate and Grounding precision. These results confirm the high quality of annotations in our OminiAbnorm-CT-14K.
A.6 The Long-Tail Distribution of OminiAbnorm-CT-14K
To mitigate the long-tail distribution in OminiAbnorm-CT-14K, we identified underrepresented organs in our annotated corpus and strategically employed GPT-4o to analyze unlabeled reports. Using the following prompt, we efficiently filtered cases containing abnormal findings related to these underrepresented organs:
This is a report of a CT scan: \$Report\$ Please help me carefully check if the report mentions any abnormal findings that belong to the following anatomical areas: \$List of Underrepresented Organs\$. If there are, please output ’YES’, otherwise output ’NO’. Do not output any other information.
These identified cases were then prioritized in our annotation pipeline. To evaluate the effectiveness of this strategy, we randomly sampled 1,000 images before and after implementation. As shown in Fig. 7, prior to our intervention, the top 20 organs (out of 82 total) accounted for 79.8% of all annotations, while the top 85 abnormality categories (out of 340 total) constituted 80.1%. Following our strategy, the underrepresented organs and rare abnormality categories received notably increased annotation coverage: the representation of underrepresented organs increased by 19.6%, while rare abnormality categories increased by 11.6%, significantly enhancing the diversity of our dataset.
Figure 7: The distribution before and after prioritizing annotation for underrepresented organs. (a) Comparison of organ distribution in annotations; (b) Comparison of abnormality category distribution in annotations.
A.7 Comparison between OminiAbnorm-CT-14K and Existing Datasets
We compare OminiAbnorm-CT-14K with existing public CT image datasets in Table 6, including those widely used for lesion segmentation, lesion detection, organ segmentation, or report generation. In contrast to the substantial limitations exhibited by these datasets, as detailed in Sec. 3.1, OminiAbnorm-CT-14Kepresents the first large-scale dataset designed for abnormality grounding and description across multi-plane and whole-body CT images.
Table 6: Comparison of key characteristics between OminiAbnorm-CT-14K and widely-used public CT imaging datasets. Note that CCC18 and DeepLesion has no category label for each annotated lesion. Even though some scans in these datasets are isotropic, they are acquired as axial-plane. Abbreviations: Rpt = Report; Cat = Abnormality Category; Img = Image; Seg. = Segmentation; Det. = Detection; Gen = Generation.
Dataset Task Anatomy Plane Rpt #Cat #Img
[…]
Lesion Seg. Axial 340 10K
OminiAbnorm-CT-14K Lesion Det. Whole Body Coronal 255 2K
Report Gen. Sagittal 223 1.6K
A.8 Ablation Study on Context Range
In clinical practice, the contextual information provided by adjacent slices is critical for the interpretation of CT images, which is also widely validated in tasks such as medical image segmentation [26] and diagnosis [64]. To investigate the impact of context range on report generation, in the visual prompted report generation task, we adjust the number of input slices and evaluate its impact on a base QWen2.5-VL-7B model. As shown in Table 7, when we increase the input slices, there is no consistent performance improvement on any metric. We speculate this is mainly because multi-modal language models, which generally encode and project each input image into tokens independently, can hardly capture the subtle spatial relationships across adjacent slices. Therefore, we only take the center slice as image input to the multi-modal language model in the subsequent training and evaluation of OminiAbnorm-CT.
Table 7: Impact of multi-slice context range on visual prompted report generation, with the base Qwen2.5-VL-7B model as baseline. Abbrevations: B = BLEU, RTSc = RaTEScore, BTSc = BERTScore, MTR = METEOR, R = Rouge, RG = RadGraph.
Input Slices B-1 B-2 B-3 RTSc BTSc MTR R-1 R-L RG
Axial (n=2193)
4 Adjacent Slices 9.39 2.98 0.45 31.12 83.52 14.90 11.51 9.01 2.46
2 Adjacent Slices 9.24 3.03 0.46 31.47 83.49 15.05 11.41 8.99 2.49
No Adjacent Slices 11.57 4.20 0.76 38.35 84.10 16.68 14.26 11.09 3.85
Coronal (n=750)
4 Adjacent Slices 10.93 3.93 0.85 32.95 83.49 15.82 13.29 9.79 3.62
2 Adjacent Slices 10.89 3.85 0.73 33.06 83.48 15.87 13.34 9.85 3.71
No Adjacent Slices 13.62 3.54 0.66 37.26 84.19 13.49 14.97 11.47 3.37
Sagittal (n=591)
4 Adjacent Slices 9.39 2.98 0.45 31.12 83.52 14.90 11.51 9.01 2.46
2 Adjacent Slices 11.48 4.16 0.94 33.86 83.32 15.70 13.64 9.97 3.75
No Adjacent Slices 13.28 3.54 0.63 35.67 83.98 13.01 14.75 11.34 2.91
A.9 Ablation Study on Different Visual Prompts
Recent studies have revealed that VLMs exhibit varying perceptual capabilities for different forms of visual prompts [8]. Therefore, in the visual-prompted report generation task, we investigated the impact of different visual prompt formats to give insight on optimal design choices in application. Specifically, for each abnormality in our dataset, we simulated four distinct visual prompt types: center cropping, ellipse, contour, and bounding box, as input and evaluated OminiAbnorm-CT’s generated reports respectively. We also take the maximum score among them for each abnormality, representing the theoretical best performance achievable by selecting the optimal visual prompt type for each specific abnormality, depending on its shape, location, and etc. As shown in Table 8, cropping the prompted area consistently underperforms across nearly all metrics (24 out of 27 total measurements), likely due to the loss of crucial contextual information outside the prompted region. The other individual visual prompting methods (ellipse, contour, and bounding box) demonstrate comparable performance. Notably, integrating multiple visual prompting techniques yields a substantial improvement across all metrics compared to the best individual prompt for each measure. This indicates significant performance variation when using different visual prompts for the same abnormality, suggesting that selecting appropriate prompt types based on abnormality morphology and location could be an effective strategy.
Table 8: Comparison of different visual prompt formats in visual prompted report generation task, with OminiAbnorm-CT fixed as baseline. Abbrevations: B = BLEU, RTSc = RaTEScore, BTSc = BERTScore, MTR = METEOR, R = Rouge, RG = RadGraph.
Prompt Type B-1 B-2 B-3 RTSc BTSc MTR R-1 R-L RG
Axial (n=2193)
Center Cropping 11.70 2.87 0.68 33.03 84.84 12.77 14.58 11.75 3.92
Ellipse 12.43 3.60 1.15 35.14 84.87 13.67 15.55 12.63 5.27
Contour 12.60 3.79 1.22 35.74 85.09 14.12 15.97 12.92 5.81
Bounding Box 12.35 3.69 1.26 35.46 84.92 13.65 15.76 12.80 5.63
Max 18.03 6.65 1.80 42.81 86.35 19.40 21.66 17.61 9.98
Coronal (n=750)
Center Cropping 10.94 2.96 0.80 33.57 84.48 11.92 14.23 11.36 3.87
Ellipse 12.49 3.79 1.13 36.09 84.56 13.44 15.94 12.43 5.16
Contour 12.61 3.73 1.02 35.50 84.71 13.45 15.75 12.25 4.87
Bounding Box 12.15 3.70 1.12 35.12 84.38 13.27 15.68 12.13 4.68
Max 18.38 7.19 2.39 42.88 86.00 19.00 21.30 16.87 9.44
Sagittal (n=591)
Center Cropping 10.98 2.73 0.39 32.23 84.20 11.98 14.45 11.01 3.05
Ellipse 11.37 3.10 0.65 34.36 84.01 12.19 14.45 10.95 3.46
Contour 11.08 2.95 0.57 34.19 84.22 12.01 14.38 11.10 3.22
Bounding Box 11.29 2.98 0.67 34.07 84.13 12.16 14.84 11.24 3.45
Max 17.29 6.48 1.82 41.77 85.62 17.45 20.43 15.87 7.51
A.10 Generation of Text Queries and Visual Prompts
We simulate four visual prompts for each annotated abnormality, mimicking how clinicians would select the most appropriate highlighting method based on the abnormality’s shape and location: (i) cropped region. We extract the minimum bounding box that completely contains the annotated lesion region, add a 50-pixel padding around this region, and crop the original image to center the abnormality; (ii) ellipse. We fit an optimal ellipse to the largest contour extracted from the lesion mask. (iii) contour. We smooth the lesion mask using Gaussian blur, detect its contours, and refine them with polygon approximation. (iv) bounding box. We identify the minimum bounding box that completely contains the annotated lesion region, and add a 10-pixel padding around.
To simulate radiologists approaching CT images with prior knowledge, we provide GPT-4o with patient information (complaints, medical history, etc) to generate text queries that inquire specific abnormality based on such preliminary information. We use the following prompt template:
Assuming a CT image has one or more abnormal findings, I will provide detailed information about them. Please help me generate prompts to test a VLM’s ability to localize and analyze specific abnormalities. Requirements: 1. Generate prompts in English. 2. Ensure accurate information. The prompt content must stem from the real abnormality information, CT image reports, and doctor’s discussion results I provided. You may use equivalents expressions in medical terminology, but do not introduce any content beyond these provided information. 3. Clear indication. Ensure each prompt refers to the corresponding abnormality without confusion with other abnormalities in the image. 4. Clear task. Each prompt should end with a clear request for the VLM to perform localization and analysis tasks. 5. Simulate realistic pre-examination clinical queries. Create prompts that reflect how clinicians approach CT images with preliminary information (such as the patient’s complaints, medical history, other test results, etc., if available) and medical knowledge (such as common abnormalities associated with the patient’s information or typically found in this imaged region), without being overly specific. Importantly, prompts should be broad enough to guide examination of suspicious areas and must not include detailed descriptions or conclusions from findings, reports and discussion results that would only be available after examining the CT image, such as specific abnormality details, exact measurements, precise locations, or definitive characteristics that could only be determined after image interpretation. Some appropriate examples: \$Some Example Queries\$ 6. Avoid data leakage. To evaluate the VLM’s ability to localize and analyze abnormalities independently, do not provide complete findings or diagnostic conclusions in the query. This prevents the VLM from bypassing the analytical process by retrieving answers directly from the prompt, while maintaining challenge authenticity. 7. Diversity. Generate at least 1 and at most 5 prompts with different perspectives for each abnormality. Each prompt should have clearly different focuses, avoiding content redundancy. The information about all the abnormalities on the CT images: \$Abnormal Findings\$ The clinical presentation of the patient corresponding to this CT image is: \$Presentation\$ The overall report for the CT image containing these abnormalities is: \$Whole Report\$ The doctors’ discussion results for the patient corresponding to this CT image are: \$Impression\$ Your response must follow this exact format: \$Output Template\$
A.11 Detailed Distribution of Abnormalities in OminiAbnorm-CT-14K
Figure 8: Distribution of OminiAbnorm-CT-14K across 82 anatomical structures and 40 major systems or organs. Darker blue indicates higher sample density. Abd. & Pelv. = Abdominal and Pelvic Cavity; Test., Epid., Scrot. = Testes, Epididymis, Scrotum; Morph. = Morphology; Pos. = Positional; Vent. & Subarach. = Ventricles and Subarachnoid.
Table 9: Detailed distribution of abnormality categories in OminiAbnorm-CT-14K.
Organ Anatomical Structure Category Axial Coronal Sagittal
Brain Cerebral parenchyma Brain parenchymal atrophy 31 2 4
Brain parenchymal edema 49 13 4
Brain parenchymal soft tissue mass 280 57 57
Brain parenchymal 52 1 3
thin-walled cystic mass
Brain parenchymal 17 0 1
thick-walled cystic mass
Brain parenchymal 198 42 31
hemorrhage or contusion
Acute infarct 88 1 0
Lacunar infarct 54 11 8
Encephalomalacia 27 6 3
Brain parenchymal 102 31 23
morphological altrenation
Other non-mass effect lesions 44 5 5
Intracranial air 14 8 10
Hyperdense lesions in brain parenchyma 72 24 44
Others 35 13 6
Ventricles and cisterns Ventricular or cisternal enlargement 122 25 25
Ventricular or cisternal soft tissue mass 49 8 6
Ventricular or cisternal cystic mass 13 1 0
Ventricular or cisternal hemorrhage 60 4 3
Others 11 2 3
Meninges (including dura mater, pia mater and arachnoid mater) Meningeal cystic mass 4 0 0
Meningeal soft tissue mass 8 9 6
Meningeal hemorrhage 137 27 10
Meningeal effusion 32 9 3
Meningeal thickening 6 0 0
Others 4 0 2
Pituitary and Sellar Region Pituitary stalk thickening 0 0 0
Pituitary stalk lateral displacement 0 0 0
Pituitary enlargement 0 0 2
Pituitary atrophy 0 2 4
Pituitary calcification 0 0 3
Pituitary or sella region soft tissue mass 20 16 27
Pituitary or sella region cysitc mass 3 5 6
Others 1 0 0
Spine Spinal Cord Spinal cord compression 3 0 2
Spinal cord soft tissue mass 12 1 6
Spinal cord hemorrhage or contusion 0 0 0
Spinal morphological alteration 0 0 0
Syringomyelia 0 0 0
Others 4 0 1
Intervertebral disc Intervertebral disc 4 2 1
morphological alteration
Intervertebral disc 0 0 0
ossification or calcification
Intervertebral disc gas 4 3 2
Intervertebral disc extrusion 0 0 0
Intervertebral disc soft tissue mass 5 1 1
Others 0 0 1
Spinal meninge (including dura mater, arachnoid mater, and pia mater) Spinal meningeal hemorrhage 0 0 0
Spinal meningeal effusion 0 0 1
Spinal meningeal thickening 0 0 0
Meningeal soft tissue mass 7 2 6
Meningeal cystic mass 0 2 0
Others 2 1 5
Table 10: (Continued) Detailed distribution of abnormality categories in OminiAbnorm-CT-14K.
Organ Anatomical Structure Category Axial Coronal Sagittal
Eye Eyeball Eyeball atrophy 4 1 2
Eyeball positioning or morphological alteration 44 7 6
Eyeball density alteration 45 20 14
Eyeball soft tissue mass 19 3 1
Eyeball wall thickening 5 3 0
morphological alteration
Complete or partial absence of eyeball structure 2 0 0
Others 2 2 0
Ocular Adnexa Orbital density changes 54 36 15
Intraorbital gas 4 4 3
(e.g., soft tissue mass, fluid accumulation)
Extraocular muscle hypertrophy or atrophy 16 43 11
Optic nerve thickening or soft tissue mass 7 3 7
Optic nerve atrophy 0 0 0
Others 3 2 3
Lacrimal gland and lacrimal sac Lacrimal gland and lacrimal sac 19 12 5
enlargement or mass
Lacrimal gland and lacrimal sac 4 2 0
calcification or fluid
Others 1 0 1
Ear External Ear External auditory canal stenosis or atresia 5 1 2
External auditory canal soft tissue mass 11 5 1
Others 1 1 0
Middle Ear Middle ear fluid or hemorrhage 5 0 0
Middle ear soft tissue mass 23 23 3
Ossicular chain destruction or deformity 7 2 0
Tympanic membrane thickening or calcification 2 2 0
Middle ear gas density change 6 0 1
Others 1 2 0
Inner Ear Inner ear congenital structural alteration 2 1 0
Inner ear bone destruction or sclerosis 11 4 0
Internal auditory canal enlargement or narrowing 2 0 0
Labyrinthine structural alteration 4 2 0
Others 1 1 0
Sinus Sinus cavity Sinus effusion 18 8 0
Sinus hemorrhage 6 3 3
Sinus soft tissue mass 98 85 33
Sinus cystic mass 8 5 1
Sinus mucosal thickening 24 25 7
Others 9 3 2
Sinus ostium Sinus obstruction or stenosis 1 4 0
Sinus widening 0 0 1
Others 1 1 0
Nasal septum Nasal septum deviation or thickening 7 6 0
Nasal septum perforation or defect 1 1 0
Others 6 2 0
Table 11: (Continued) Detailed distribution of abnormality categories in OminiAbnorm-CT-14K.
Organ Anatomical Structure Category Axial Coronal Sagittal
Pharynx Pharynx (including nasopharynx, oropharynx, and hypopharynx) Pharyngeal narrowing or obstruction 4 2 1
Pharyngeal soft tissue mass 46 14 10
Pharyngeal cystic mass 10 2 4
Pharyngeal wall thickening 4 3 0
Pharyngeal foreign body 0 0 0
Others 3 4 1
Larynx Laryngeal narrowing or obstruction 3 2 0
Vocal cord asymmetry 2 0 0
Vocal cord soft tissue mass 6 1 1
Laryngeal cartilage calcification 0 0 0
Laryngeal cystic mass 1 1 0
Laryngeal foreign body 1 0 0
Others 1 0 0
Pharyngeal space Pharyngeal space soft tissue mass 46 23 16
Pharyngeal space cystic mass 13 6 2
Pharyngeal emphysema 7 0 0
Others 5 7 1
Parotid gland Parotid gland Parotid gland enlargement 20 2 0
Parotid gland atrophy 4 0 0
Parotid gland soft tissue mass 33 2 2
Parotid gland cystic mass 16 0 0
Parotid gland calcification or stone 15 0 2
Others 10 0 0
Thyroid gland Thyroid gland Thyroid enlargement 12 2 2
Thyroid atrophy 1 0 0
Thyroid soft tissue mass 41 4 0
Thyroid cystic mass 3 1 0
Thyroid calcification 1 0 0
Ectopic thyroid gland 5 1 0
Others 6 0 1
Trachea Tracheal lumen Tracheal stenosis or obstruction 12 4 7
Tracheal dilatation 11 5 2
Tracheal soft tissue mass 12 6 1
Tracheal wall thickening 9 0 0
Tracheal wall calcification 3 1 4
Tracheal wall defect 3 0 2
Others 1 0 2
Lung pulmonary parenchyma Atelectasis 106 10 3
Incomplete lung expansion 52 3 1
Pulmonary consolidation 203 24 2
Pulmonary ground-glass opacities 343 52 5
Pulmonary emphysema 45 10 2
Pulmonary solitary nodule or mass 226 46 10
Pulmonary diffusely 285 76 3
distributed multiple nodules
Pulmonary parenchymal fibrosis 42 12 0
Pulmonary thin-walled cavitation 48 6 1
Pulmonary thick-walled cavities 44 1 0
Pulmonary cystic mass 53 14 1
Others 14 6 1
Bronchi Bronchiectasis 97 19 6
Bronchial wall thickening, 15 11 5
stenosis, or occlusion
Bronchial foreign body 4 1 0
Others 3 0 1
Table 12: (Continued) Detailed distribution of abnormality categories in OminiAbnorm-CT-14K.
Organ Anatomical Structure Category Axial Coronal Sagittal
Lung Lung interstitial Lung interstitial fibrosis and thickening 78 10 0
Honeycomb lung 34 0 0
Others 2 0 0
Pleura Pleural thickening 39 10 1
Pleural effusion 268 27 3
Pneumothorax 36 2 0
Pleural calcification 24 0 1
Pleural soft tissue mass 50 8 0
Others 8 1 1
Mediastinum Mediastinal soft tissue Mediastinal shift 29 4 0
Mediastinal soft tissue mass 200 31 15
Mediastinal cystic mass 19 0 2
Mediastinal hemorrhage 15 0 0
Mediastinal emphysema 58 7 1
Others 10 0 0
Diaphragm Diaphragm Diaphragmatic hernia 19 16 12
Diaphragmatic elevation 5 4 0
Diaphragmatic soft tissue mass 7 2 0
Others 0 2 0
Thymus Thymic parenchyma Thymic enlargement 1 0 0
Thymic atrophy 0 0 0
Thymic soft tissue mass 6 0 0
Thymic cystic mass 0 0 0
Thymic calcification 4 0 0
Others 1 0 0
Oral cavity Oral Soft Tissue Oral soft tissue mass 8 0 0
Oral cystic mass 3 0 0
Oral soft tissue calcification 0 0 0
Others 0 0 0
Teeth and alveolar bone Dental developmental anomalies 1 0 0
Dental positional anomalies 3 4 5
Dental calcification or caries 3 0 0
Alveolar bone resorption or hyperplasia 3 0 0
Alveolar soft tissue mass 8 2 0
Alveolar cystic mass 8 1 1
Others 0 1 0
Heart Cardiac chambers (atrium or ventricle) Cardiac chamber enlargement 41 7 3
Cardiac chamber mass 28 7 2
Others 17 1 0
Myocardium Myocardial hypertrophy 4 1 1
Myocardial thinning 3 0 0
Myocardial calcification 4 1 0
Myocardial density alteration 6 0 0
Others 3 3 0
Heart valve Valvular calcification 3 0 0
Others 3 0 0
Pericardium Pericardial thickening 9 0 2
Pericardial calcification 6 1 0
Pericardial effusion 54 7 2
Pericardial hemorrhage 2 0 0
Pericardial emphysema 6 1 0
Others 4 0 0
Table 13: (Continued) Detailed distribution of abnormality categories in OminiAbnorm-CT-14K.
Organ Anatomical Structure Category Axial Coronal Sagittal
Heart Coronary arteries Coronary artery myocardial bridge 1 0 0
Coronary artery dilation 0 0 0
Others 3 1 0
Vascular structure Artery Arterial widening 66 24 9
Aneurysm 138 44 39
Atherosclerotic plaque 39 13 4
of arterial wall
Arterial wall ulcer 8 2 2
Arterial dissection or 22 15 8
intramural hematoma
Arterial stenosis 24 8 10
Arterial occlusion 55 9 0
Arterial filling defect 80 23 8
Arterial contour abnormality 51 23 23
Arteriovenous fistula 16 7 1
Arterial wall inflammatory exudate 4 0 0
Others 39 9 7
Vein Venous dilation 41 4 6
Varicosity 23 11 9
Venous wall inflammation 0 0 0
Venous stenosis 7 1 1
Venous occlusion 5 2 1
Venous filling defect 48 12 3
Venous morphological abnormality 29 6 3
Venous wall inflammatory exudate 2 1 0
Others 13 8 0
Capillary Capillary dilation 0 0 0
Capillary malformation proliferation 1 0 0
Others 0 0 0
Gastrointestinal tract (including esophagus, stomach, and intestines) Gastrointestinal lumen Gastrointestinal dilatation 265 212 112
Gastrointestinal narrowing 20 5 6
Gastrointestinal foreign body 2 6 0
Gastrointestinal air-fluid level 31 8 5
Gastrointestinal luminal 17 6 4
contents abnormal density
Others 8 3 1
Gastrointestinal tract wall Gastrointestinal wall thickening 398 212 120
Gastrointestinal wall mass 103 56 26
Gastrointestinal wall 43 12 4
rupture and perforation
Gastrointestinal wall ulceration 0 0 0
Others 7 9 5
Gastrointestinal positioning abnormality Gastrointestinal herniation 117 67 30
or deformity
Others 5 1 1
Gastrointestinal morphological abnormalities Gastrointestinal diverticulum 53 35 15
Gastrointestinal malrotation and volvulus 17 17 3
Gastrointestinal annular or 28 22 10
concentric abnormality
Others 2 1 2
Mesentery Mesenteric volvulus 3 4 2
Mesenteric edema 10 3 3
Mesenteric panniculitis 26 6 6
Mesenteric soft tissue mass 28 6 9
Others 0 4 2
Table 14: (Continued) Detailed distribution of abnormality categories in OminiAbnorm-CT-14K.
Organ Anatomical Structure Category Axial Coronal Sagittal
Liver Hepatic parenchyma Hepatic parenchymal morphological alteration 99 34 11
Hepatic parenchymal hyperdensity 18 2 0
Hepatic parenchymal hypodensity 71 25 9
Hepatic parenchymal soft tissue mass 439 58 20
Hepatic parenchymal thin-walled cystic mass 124 23 10
Hepatic parenchymal thick-walled cystic mass 21 6 4
Liver contusion or hemorrhage 10 3 0
Intrahepatic bile duct emphysema 13 1 0
Intrahepatic bile duct fluid accumulation 0 0 0
Intrahepatic biliary stones or calcification 7 0 0
Others 16 4 1
Gallbladder Gallbladder lumen Gallbladder distension 39 10 1
Gallbladder atrophy or shrinkage 4 1 0
Gallbladder stone 87 22 2
Gallbladder contents density change 32 5 0
Others 3 1 0
Gallbladder wall Gallbladder wall thickening 53 11 1
Gallbladder wall calcification 9 1 0
Gallbladder wall mass 16 2 1
Gallbladder wall rupture 11 1 0
Others 3 1 1
Gallbladder morphology Gallbladder position alteration 1 0 0
Gallbladder congenital morphological variation 0 0 0
Others 1 1 0
Extrahepatic bile ducts Extrahepatic bile duct dilation 26 8 0
Extrahepatic bile duct wall thickening 2 0 0
Extrahepatic bile duct soft tissue mass 1 0 1
Extrahepatic bile duct cystic mass 1 1 0
Extrahepatic bile duct stenosis or obstruction 2 0 0
Extrahepatic bile duct content density alteration 6 1 0
Extrahepatic bile duct stone 15 5 0
Extrahepatic bile duct injury or rupture 0 0 0
Others 1 0 0
Pancreas Pancreatic parenchyma Pancreatic parenchymal soft tissue mass 137 8 4
Pancreatic cystic mass 69 7 4
Pancreatic calcification 20 1 0
Pancreatic enlargement 60 1 1
Pancreatic atrophy 8 1 1
Others 16 0 0
Pancreatic duct Pancreatic ductal dilatation 13 3 0
Pancreatic ductal stone 3 0 0
Others 0 0 0
Pancreatic morphology Pancreatic congenital anomaly 5 0 0
Pancreatic positional displacement 1 0 0
Others 1 0 0
Spleen Splenic parenchyma Splenic parenchymal calcification 15 0 1
Splenic parenchymal soft tissue mass 52 11 3
Splenic parenchymal cystic mass 27 10 0
Splenic parenchymal infarct 19 1 0
Splenic parenchymal rupture 12 1 0
Others 18 2 0
Spleen morphology Spleen enlargement 102 22 5
Others 17 1 0
Table 15: (Continued) Detailed distribution of abnormality categories in OminiAbnorm-CT-14K.
Organ Anatomical Structure Category Axial Coronal Sagittal
Abdominopelvic peritoneum Abdominopelvic peritoneum Peritoneal inflammatory exudate 75 19 10
Peritoneal thickening 19 3 1
Peritoneal calcification 5 1 0
Peritoneal soft tissue mass 364 135 90
Peritoneal cystic mass 102 34 22
Abdominopelvic fluid 280 96 43
Abdominopelvic hemorrhage 68 3 1
Abdominopelvic free air 61 17 10
Retroperitoneal fibrosis 1 0 0
Peritoneal or retroperitoneal 17 18 4
lymph enlargement
Extravasation of gastrointestinal content 0 1 0
Abdominopelvic contrast agent leakage 22 13 4
Others 12 12 2
Kidney Renal Parenchyma Renal parenchymal soft tissue mass 150 81 29
Renal parenchymal cystic mass 239 68 21
Renal parenchymal calcification 9 9 2
Renal parenchymal or 11 3 2
subcapsular hemorrhage
Renal infarct 8 3 0
Others 27 8 1
Renal pelvis and ureter Hydronephrosis 81 63 3
Ureteral dilatation 35 29 4
Ureteral stricture or obstruction 1 1 0
Ureteric stone 59 60 19
Double renal pelvis and/or 5 3 1
double ureter anomaly
Renal pelvis soft tissue mass 13 4 0
Renal pelvic cystic mass 1 3 0
Others 4 5 2
Renal morphology and position abnormalities Renal morphological anomaly 33 20 1
Renal enlargement 56 16 1
Renal atrophy 23 6 2
Ectopic or transplanted kidney 17 13 5
Others 3 2 0
Bladder Bladder cavity Bladder distention 4 0 1
Bladder stone 24 5 3
Bladder content density change 13 2 1
Others 4 3 0
Bladder wall Bladder wall diffuse thickening 17 5 6
Bladder wall calcification 1 3 1
Bladder wall focal thickening 69 7 3
or soft tissue mass
Bladder wall defect or fistula 4 0 1
Bladder wall emphysema 9 0 1
Bladder wall diverticulum 31 2 2
Others 1 0 0
Bladder morphology Bladder position displacement 9 2 2
Bladder morphological anomaly 2 0 0
Others 1 0 2
Table 16: (Continued) Detailed distribution of abnormality categories in OminiAbnorm-CT-14K.
Organ Anatomical Structure Category Axial Coronal Sagittal
Adrenal gland Adrenal gland Adrenal soft tissue mass 122 17 1
Adrenal cystic mass 32 0 0
Adrenal calcification 0 3 0
Adrenal gland thickening 16 0 0
Adrenal atrophy 0 0 0
Others 15 0 0
Prostate Prostate Prostate enlargement 18 12 14
Prostate atrophy 0 0 0
Prostatic cystic mass 2 0 0
Prostatic soft tissue density anomaly 2 0 1
Prostatic calcification 2 1 2
Prostatic hemarrhage 0 0 0
Others 0 0 0
Seminal vesicle Seminal vesicle Seminal vesicle soft tissue mass 1 0 0
Seminal vesicle calcification 0 0 1
Seminal vesicle cystic mass 3 0 0
Others 0 0 0
Testes, epididymis, and scrotum Testis Testicular enlargement 1 0 0
Testicular atrophy 0 0 0
Testicular soft tissue mass 1 0 0
Testicular calcification 0 0 0
Testicular cystic mass 0 0 0
Testicular torsion 0 0 0
Testicular hemorrhage and rupture 0 0 0
Others 0 0 0
Scrotum Scrotal effusion 1 0 0
Scrotal hematoma 0 0 0
Scrotal soft tissue mass 1 0 0
Scrotal wall thickening 0 0 0
Others 2 0 0
Epididymis Epididymis enlargement 0 0 0
Epididymal soft tissue mass 0 0 0
Epididymal calcification 0 0 0
Epididymal cystic mass 0 0 0
Epididymal thickening 0 0 0
Others 0 0 0
Penis Penis Penile morphological anomaly 0 0 0
Penile soft tissue mass 0 0 0
Penile calcification 0 0 0
Urethral calculi or foreign body 6 2 1
Urethral Stricture 0 0 0
Urethral dilation 1 0 0
Others 4 0 0
Uterus Uterus Uterine morphological anomaly 7 1 2
Uterine enlargement 15 3 6
Uterine soft tissue mass 50 4 7
Uterine calcification 7 0 2
Uterine cavity effusion 2 0 1
Uterine cavity hemorrhage 2 0 0
Uterine cystic mass 4 0 0
Others 4 1 1
Table 17: (Continued) Detailed distribution of abnormality categories in OminiAbnorm-CT-14K.
Organ Anatomical Structure Category Axial Coronal Sagittal
Fallopian tube Fallopian tube Fallopian tube thickening 3 0 0
Fallopian tube cystic mass 2 0 1
Fallopian tube soft tissue mass 12 4 2
Fallopian tube effusion 2 0 0
Fallopian tube hemorrhage 0 0 0
Fallopian tube calcification 0 0 0
Others 0 0 0
Ovary Ovary Ovarian enlargement 9 3 0
Ovarian Atrophy 0 0 0
Ovarian cystic mass 85 12 6
Ovarian soft tissue mass 96 6 3
Ovarian calcification 3 1 0
Ovarian torsion 9 0 0
Others 1 1 0
Vagina and vulva Vagina and vulva Vaginal soft tissue mass 0 0 0
Vaginal cystic mass 9 0 0
Vaginal hemorrhage 1 0 0
Vaginal emphysema 3 0 4
Vaginal anatomical anomaly 3 0 1
Others 6 0 0
Breast Breast gland Breast gland enlargement 2 0 0
Breast gland atrophy 0 0 0
Breast gland soft tissue mass 36 9 2
Breast gland calcification 0 0 0
Breast gland cystic mass 0 0 0
Others 6 0 0
Breast duct Breast duct dilation 0 0 0
Breast duct calcification 0 0 0
Others 0 0 0
Nipple Nipple retraction 0 0 0
Nipple calcification 0 0 0
Others 0 0 0
Areola Areola thickening 1 0 0
Others 0 0 0
Skeletal system Skeletal system Osteoporosis 9 1 8
Osteomalacia 2 0 1
Bone destruction or soft tissue mass 219 124 117
Bone cystic mass 19 32 17
Osseous sclerosis 179 80 121
Osteonecrosis 64 13 24
Bone fracture 308 172 247
Periosteal reaction 23 9 5
Periosteal thickening 8 5 1
Bone callus and post-fracture healing 10 8 2
Scar of fracture fixation removal 0 0 0
Bone deformation 67 44 73
Skeletal asymmetry 10 2 1
Cartilage calcification 4 4 8
Chondral calcification 5 0 1
Others 30 50 26
Table 18: (Continued) Detailed distribution of abnormality categories in OminiAbnorm-CT-14K.
Organ Anatomical Structure Category Axial Coronal Sagittal
Joint Joint Joint space narrowing 6 12 3
Joint space widening 3 0 3
Joint cartilage degradation 1 0 2
Joint cartilage calcification 3 1 1
Joint capsule thickening 0 0 0
Intra-articular effusion 13 3 2
Intra-articular hemorrhage 4 0 0
Intra-articular gas 1 0 4
Joint periarticular 2 0 0
soft tissue swelling
Irregular articular surfaces 12 5 2
Joint subluxation or dislocation 30 14 22
Intra-articular loose body 8 1 2
Others 5 12 2
Muscle Muscle Muscle swelling 8 5 1
Muscle atrophy 13 1 0
Muscular soft tissue mass 40 9 2
Muscular cystic mass 11 13 3
Muscular hemorrhage 7 0 1
Muscle calcification 3 0 0
Muscle open injury and tear 0 0 0
Tendon calcification 19 3 13
Tendon tear or rupture 1 0 1
Others 7 4 1
Skin and subcutaneous fat Skin and subcutaneous fat Subcutaneous soft tissue mass 141 47 28
Subcutaneous edema 18 6 2
Subcutaneous effusion 26 3 1
Subcutaneous inflammatory 10 2 1
exudate
Subcutaneous swelling 60 10 15
Subcutaneous open wound 15 5 2
and laceration
Subcutaneous calcification 12 0 0
Subcutaneous fat necrosis 2 0 0
Abdominal wall hernia 48 3 6
Others 33 5 9
Agenesis or ectopia Agenesis or ectopia Congenital developmental 78 35 14
anomaly
Postoperative changes 42 16 5
Situs inversus 10 7 1
Others 2 0 0
Implantation of artificial object Implantation of artificial object Implantation of artificial object 179 91 53
Others 17 3 8

### v2

@@
 Automated interpretation of CT images—particularly localizing and describing abnormal findings across multi-plane and whole-body scans—remains a significant challenge in clinical radiology.
-This work aims to address this challenge through four key contributions: (i) On taxonomy, we collaborate with senior radiologists to propose a comprehensive hierarchical classification system, with 404 representative abnormal findings across all body regions; (ii) On data, we contribute a dataset containing over 14.5K CT images from multiple planes and all human body regions, and meticulously provide grounding annotations for over 19K abnormalities, each linked to the detailed description and cast into the taxonomy; (iii) On model development, we propose OminiAbnorm-CT, which can automatically ground and describe abnormal findings on multi-plane and whole-body CT images based on text queries, while also allowing flexible interaction through visual prompts; (iv) On benchmarks, we establish three representative evaluation tasks based on real clinical scenarios.
-Through extensive experiments, we show that OminiAbnorm-CT can significantly outperform existing methods on all the tasks and metrics.
+This work aims to address this challenge through four key contributions: (i) On taxonomy, we collaborate with senior radiologists to propose a comprehensive hierarchical classification system, with 404 representative abnormal findings across all body regions; (ii) On data, we contribute a dataset containing over 14.5K CT images from multiple planes and all human body regions, and meticulously provide grounding annotations for over 19K abnormalities, each linked to the detailed description and cast into the taxonomy; (iii) On model development, we propose OmniAbnorm-CT, which can automatically ground and describe abnormal findings on multi-plane and whole-body CT images based on text queries, while also allowing flexible interaction through visual prompts; (iv) On evaluation, we establish three representative tasks based on real clinical scenarios, and introduce a clinically grounded metric to assess abnormality descriptions.
+Through extensive experiments, we show that OmniAbnorm-CT can significantly outperform existing methods in both internal and external validations, and across all the tasks.
+Figure 1: This work introduces OmniAbnorm-CT-14K, the first large-scale dataset for grounding and describing abnormal findings in CT images.
+Left: OmniAbnorm-CT-14K contains 14.5K multi-plane, whole-body CT images, covering 349 representative abnormal findings across 82 anatomical structures and 40 major systems or organs.
+Right: Distribution of the dataset across anatomical structures and major systems or organs, with darker blue indicating higher sample density.
 1 Introduction
-Computed Tomography (CT) imaging has become a cornerstone of modern medicine, with over 3 billion scans performed annually around the world.
-The interpretation of these scans—particularly the generation of radiology reports—plays a crucial role in guiding clinical decision-making.
-However, systematically identifying and characterizing abnormal findings in whole-body CT images remains a cognitively demanding and time-consuming task for radiologists rao2025multimodal .
-In the recent literature, artificial intelligence has started to make significant progress in CT image interpretation, particularly along two directions: first, AI models have been developed for organ-level segmentation Totalsegmentator ; jaus2023towards , that enable precise anatomical labeling and organ-grounded report generation zhang2024radgenome ; chen2025large .
+Computed Tomography (CT) imaging has become a cornerstone of modern medicine, with over 450 million scans performed annually worldwide.
+The interpretation of these scans, particularly the generation of radiology reports, plays a crucial role in clinical decision-making.
+However, systematically identifying and characterizing abnormal findings in whole-body CT images remains a cognitively demanding and time-consuming task for radiologists [rao2025multimodal].
+Advances in artificial intelligence (AI) have shown promising progress in CT image interpretation, primarily along two directions: first, models developed for organ segmentation [Totalsegmentator, jaus2023towards] enable precise anatomical labeling and organ-grounded report generation [zhang2024radgenome, chen2025large].
 Yet a critical gap remains: segmenting organs alone is insufficient for clinical utility.
-What ultimately matters to clinicians are abnormal findings—lesions and any anomalies that inform diagnosis and treatment; second, models for automatic report generation have been trained on the recently introduced datasets with paired CT scans and corresponding radiology reports, describing clinical findings and impressions chen2024bimcv ; hamamci2024foundation .
-However, these datasets are typically limited to specific anatomical regions—most commonly chest CT—and often lack explicit visual grounding, which reduces explainability and increases the risk of hallucinated content.
-To address the above-mentioned challenges, we aim to explore the potential to develop a system that automatically detects, localizes, and describes all abnormal findings across multi-plane, whole-body CT images.
-This paradigm shift moves beyond regional reporting and anatomical segmentation, aligning AI interpretation more closely with the practical demands of radiological diagnosis.
-Figure 1: OminiAbnorm-CT-14K is the first large-scale dataset for abnormal findings grounding and description across multi-plane and whole-body CT images.
-It covers 349 representative abnormal findings from 82 anatomical structures and 40 major systems or organs.
-One crucial barrier to developing such a system is the absence of a unified, clinically meaningful taxonomy of abnormalities, which presents a major challenge for defining task scope and establishing continuous benchmarks.
-In this paper, we collaborated with 7 radiologists from 3 centers (each with 10-16 years of experience) to construct a detailed hierarchical taxonomy covering 404 representative abnormal findings, that are organized across 40 major anatomical regions and 82 sub-regions.
-Building on this foundation, we introduce OminiAbnorm-CT-14K-the first large-scale dataset designed for abnormality grounding and description on multi-plane whole-body CT imaging.
-It comprises 14.5K CT images from Radiopedia radiopaedia , covering axial, coronal, and sagittal planes and diverse anatomical regions.
-All images and the paired reports have been rigorously reviewed by experienced radiologists.
-We invite 4 radiologists (with at least 7 years’ experience) to manually annotate around 19K abnormal findings on the images in the format of either bounding box or segmentation masks.
+What ultimately matters to clinicians are abnormalities, lesions, and any anomalies that inform diagnosis and treatment; second, models for automatic report generation have been trained on the recent datasets with paired CT scans and radiology reports, describing findings and impressions [chen2024bimcv, hamamci2024developing].
+However, these datasets are typically limited to specific anatomical regions, e.g., chest CT, and often lack explicit visual grounding, which reduces explainability and increases the risk of hallucinations.
+To address these challenges, we aim to develop a system that detects, localizes, and describes all abnormal findings across multi-plane, whole-body CT images automatically.
+This paradigm shift moves beyond regional reporting and anatomical segmentation, aligning AI interpretation more closely with practical radiological demands.
+One crucial barrier to developing such a system is the absence of a unified, clinically meaningful taxonomy of abnormalities, which hinders the definition of task scope and continuous benchmarks.
+In this paper, we collaborated with 7 senior radiologists from 3 centers (each with 10-16 years of experience) to construct a hierarchical taxonomy covering 404 representative abnormal findings, organized across 40 major anatomical regions and 82 sub-regions.
+Based on the taxonomy, we introduce OmniAbnorm-CT-14K, the first large-scale dataset designed for abnormality grounding and description on multi-plane whole-body CT imaging.
+It comprises 14.5K CT images from Radiopedia [radiopaedia], covering axial, coronal, and sagittal planes and diverse anatomical regions.
+All images and the paired reports are contributed and rigorously reviewed by worldwide radiologists.
+We invite 4 board-certified radiologists to manually annotate around 19K abnormal findings on the images in the format of either bounding box or segmentation masks.
 All regional annotations are linked to the corresponding report descriptions, and further cast into our proposed taxonomy.
 These annotations cover 349 out of the 404 representative abnormal findings (86%) defined in our taxonomy across whole-body CT images.
-This dataset addresses a critical gap in the field, by providing high-quality, clinically relevant annotations at scale—overcoming limitations in existing datasets caused by privacy constraints and the high cost of expert annotation.
-All our annotations will be released to the community, while the images are accessible for research purposes after applying and paying the required fee to Radiopaedia [^1] https://radiopaedia.org/licence#data-sets.
-Leveraging OminiAbnorm-CT-14K, we develop OminiAbnorm-CT, a novel system for grounded CT image interpretation.
-By bridging a multi-modal language model with a specialized segmentation module in a bidirectional workflow, it enables dynamic invocation of the segmentation module during the text generation, and grounded abnormality interpretation based on the segmentation results.
-In contrast to prior approaches—which often lack grounding abilities, generalizability across planes and body regions, or flexibility in prompt handling—OminiAbnorm-CT enables automatic localization and description of abnormalities across multi-plane, whole-body CT scans.
-More importantly, OminiAbnorm-CT is designed to robustly interpret diverse textual references and visual prompts, enabling interactive and clinically adaptive use cases aligned with radiologists’ workflows.
-For a comprehensive evaluation, we define three clinically relevant tasks: (i) grounded report generation, (ii) text-guided grounded report generation, and (iii) visual-prompted report generation.
-Experiment results show that OminiAbnorm-CT significantly outperforms existing baselines across all tasks, marking a substantial step toward explainable, abnormality-centric CT imaging interpretation.
-In summary, our contributions to advance abnormality-centric grounded CT image interpretation can be summarized as: (i) We propose a comprehensive taxonomy for abnormalities on CT images, presented in Section 2; (ii) We build OminiAbnorm-CT-14K, the first dataset for abnormality grounding and description across multi-plane whole-body CT images, detailed in Section 3; (iii) We develop a novel system OminiAbnorm-CT, for automatically localizing and reporting abnormalities, illustrated in detail in Section 4; (iv) We set up a comprehensive benchmark with three clinically representative tasks, and demonstrate the superiority of OminiAbnorm-CT in Section 5.
+By providing high-quality, clinically relevant annotations at scale, this dataset addresses the critical data limitations in the field, caused by privacy constraints or high annotation cost.
+All our annotations will be released to the community, while the images can be accessed for research purposes after application [^1] https://radiopaedia.org/licence#data-sets.
+Leveraging this dataset, we develop OmniAbnorm-CT, a novel system for grounded CT image interpretation.
+By bridging a vision language model (VLM) with the specialized segmentation module in a bidirectional workflow, it enables the VLM to dynamiclly invoke the segmentation module during generation, and generate abnormality interpretation grounded on the segmentation results.
+In contrast to prior approaches, which often lack grounding abilities, generalizability across planes and body regions, or flexibility in prompt handling, OmniAbnorm-CT enables automatic localization and description of abnormalities across multi-plane, whole-body CT images.
+Moreover, OmniAbnorm-CT is designed to robustly interpret diverse textual and visual prompts, enabling interactive and adaptive use cases aligned with radiologists’ clinical workflows.
+For a comprehensive evaluation, we first define three clinically relevant tasks: (i) grounded report generation to detect, localize, and describe all findings; (ii) text-guided grounded report generation to detect, localize, and describe abnormalities based on textual queries about specific clinical concerns, and (iii) visual-prompted report generation to describe a particular abnormality indicated by a visual prompt.
+Then, we propose a clinically grounded metric for report evaluation, named AbnormRubric.
+In contrast to existing metrics relying on simple matching [bleu, rouge, banerjee2005meteor] or overall semantic similarity [zhangbertscore, zhao2024ratescore, delbrouck2024radgraph], AbnormRubric customizes the rubric for each abnormality category, ensure alignment with clinical judgement.
+Implemented using a large language model (LLM) as a judge, it explicitly quantifies recall and precision in abnormality detection, along with the descriptive accuracy of abnormalitis across four key radiological attributions: location, morphology, density, and size.
+Experiment results show that OmniAbnorm-CT significantly outperforms existing baselines across all tasks, marking a substantial step toward explainable, abnormality-centric CT imaging interpretation.
+In summary, our contributions to advance abnormality-centric grounded CT image interpretation are: (i) A comprehensive taxonomy for abnormalities on CT images, presented in Section 3; (ii) OmniAbnorm-CT-14K, the first dataset for abnormality grounding and description across multi-plane whole-body CT images, detailed in Section 4; (iii) OmniAbnorm-CT, a system for automatically localizing and reporting abnormalities, illustrated in Section 5; (iv) A benchmark with three representative tasks and a clinically grounded metrics for report evaluation, demonstrating the superiority of OmniAbnorm-CT in Section 6.
+2 Related Work
+Medical image segmentation aims to delineate clinically meaningful regions on medical images.
+For radiology, most datasets focus on annotations for important organs and anatomical structures [Totalsegmentator, abdomenatlas, FLARE, AMOS22] or lesions [heller2021kidney, MSD, BraTS2023GLI, de2024uls23, yan2018deep].
+The past decade witnessed both remarkable success in specialist segmentation models [nnUNET, UNet, zhou2023nnformer, zhou2019unet++] and a recent paradigm shift towards building generalist models [MedSAM, zhao2024foundation, zhao2025large, du2024segvol].
+However, few works contribute to the grounding of abnormal image findings, a broader spectrum including abnormal organs, potential lesions, and more generally, any anomalies meaningful for clinical decision-making.
+A concurrent work [baharoon2025rexgroundingct] provide segmentation for 14 types of abnormalities, while limited to chest CT.
+Medical image report generation aims to faithfully interpret the medical images with formal text, which is indispensable for numerous clinic procedures [rao2025multimodal].
+Existing datasets with image-report pairs are limited to specific modalities and regions, such as chest X-ray [johnson2019mimic], brain MRI [lei2024autorg] and chest CT [hamamci2024developing, chen2024bimcv], overlooking the demands for multi-plane whole-body CT report generation.
+Recent progress [zhang2024radgenome, chen2025large, lei2024autorg, chen2024vision] on grounded report generation is limited to organ-level segmentation and description, failing to localize and interpret fine-grained abnormality.
+Large visual-language models (LVLM) for medicine are built on large-scale multi-modal medical data, demonstrating exceptional performance and generalization capabilities across diverse tasks [wu2023towards, moor2023med, zhang2024generalist, li2023llava, meddr].
+Despite their potential application to multi-plane whole-body CT images, most existing approaches fail to provide grounded evidence during generation simultaneously, leading to poor explainability and hallucinations that are more difficult to detect.
 3 Taxonomy for Abnormal Findings
-Given the diversity of abnormal findings, their complex relationships and synonyms, we first construct a comprehensive taxonomy system, to categorize all the abnormalities into representative classes.
+Given the diversity, complex relationship, and synonyms of abnormalities, we first construct a taxonomy system to categorize all the abnormalities into representative classes.
 Our taxonomy is guided by four fundamental principles to ensure scientific rigor, systematic organization, and clinical utility: (i) comprehensive coverage.
 It should include all the clinically significant abnormalities, encompassing both lesions and the abnormal changes on anatomical structures; (ii) image-based definition.
-Each abnormal finding is precisely characterized by its observable features on the radiology scan, such as physical properties and morphological characteristics, that form the essential foundation for diagnosis; (iii) hierarchical organization.
-Findings are grouped by clinically important anatomical structures and major systems or organs, aligning with standard image interpretation protocols.
+Each abnormal finding is precisely characterized by its observable features on the radiology image, such as physical properties and morphological characteristics that form the essential foundation for diagnosis; (iii) hierarchical organization.
+Findings are grouped by anatomical structures and major systems or organs, aligning with standard image interpretation protocols.
 While cross-regional structures (like skeletal systems) remain independent groups for conciseness; (iv) modular design.
 Each category represents a fundamental, independent finding that can be combined with others to express complex and multifaceted abnormalities.
-The taxonomy is developed through the following three stages, involving 7 radiologists with 10 to 16 years of experience from three clinical centers, (i) we establish an anatomical hierarchy framework based on standard atlases and classic textbooks for human anatomy gray2008gray ; (ii) the radiologists then systematically catalog the abnormal findings under each anatomical structure, based on their clinical experience and several authoritative textbooks adam2020grainger ; haaga2016ct ; Mandell2013CoreRadiology on medical imaging and radiological pathology; lastly, (iii) the radiologists conduct rigorous cross-validation and discussion to establish precise consensus on the definitions of controversial or ambiguous abnormalities.
-We finally identified 404 representative abnormalities across 82 anatomies throughout the human body, organized within 40 major organs and systems. The details can be found in Appendix Section A.11.
-3 Dataset Curation
-In this section, we present OminiAbnorm-CT-14K, which establishes a data foundation for abnormality grounding and interpretation across whole-body CT images.
-The dataset construction pipeline is illustrated in Figure 2.
-Here, we start with an overview of the limitations of existing datasets in Section 3.1, which motivate the dataset construction.
-We then describe the data source in Section 3.2, detail the annotation procedure in Section 3.3, outline the extension to instruction-style data in Section 3.4, and conclude with a summary in Section 3.5.
-3.1 Limitation in Existing Datasets
-We analyze several widely used public datasets relevant to our task (summarized in Appendix Table 6) and identify the following limitations: (i) limited annotation coverage.
-Most datasets provide grounding annotations for only a small subset of lesion types.
-Despite efforts to map existing labels to our proposed taxonomy, the majority of abnormal finding categories remain unannotated; (ii) restricted imaging scope.
-Existing datasets are primarily limited to axial-plane CT images or focus on specific anatomical regions, overlooking the growing clinical need for interpretation across multi-plane, whole-body CT scans; (iii) insufficient image-report alignment.
-Many datasets lack explicit associations between localized abnormalities in the images and their corresponding descriptions in radiology reports, hindering the development of grounded, explainable systems.
-These limitations make existing datasets inadequate for benchmarking or developing models capable of automatic detection, localization, and description of abnormal findings in multi-plane, whole-body CT imaging.
+The taxonomy is developed by 7 senior radiologists with at least 10-year experience from 3 centers, following three stages: (i) an anatomical hierarchy framework is establish based on standard atlases and authoritative human anatomy textbooks [gray2008gray]; (ii) abnormal findings are systematically cataloged under each anatomical structure, based on the clinical experience and authoritative textbooks [adam2020grainger, haaga2016ct, Mandell2013CoreRadiology] on radiological imaging and pathology; (iii) the radiologists conduct rigorous cross-validation and discussion to establish precise consensus on the definitions of controversial or ambiguous abnormalities.
+We finally identified 404 representative abnormalities across 82 anatomies throughout the human body, organized within 40 major organs or systems.
+4 OmniAbnorm-CT-14K
+Figure 2: Data curation overview.
+(a) The image-report pairs are collected from an open-sourced and expert-checked website; (b) and (c) Radiologists provide grounding annotation on any abnormalities, link to their text description in reports, and categorize into the taxonomy devised by senior radiologists.
+The annotation is further extended to instruction data with simulated visual prompts and text queries.
 4.1 Data Source
-Following prior works wu2023towards ; zheng2024large ; wu2024mrgen , we collect raw data from Radiopaedia radiopaedia , a publicly accessible platform of peer-reviewed clinical cases contributed by clinicians worldwide, with appropriate privacy safeguards.
+Following prior works [wu2023towards, zheng2024large, wu2024mrgen], we collect raw data from Radiopaedia [radiopaedia], a publicly accessible platform of peer-reviewed clinical cases contributed by clinicians worldwide, with appropriate privacy safeguards.
 After application, we have obtained proper permission for all data used in this study.
-Shown in Figure 2(a), each case includes radiology scans across various planes and anatomical regions, along with patient background information and detailed clinical reports describing findings and diagnoses.
+Shown in Figure 2(a), each case includes radiology images across various planes and anatomical regions, along with patient background information and clinical reports describing findings and diagnoses.
 In total, we collected 46,721 CT images paired with 14,920 clinical reports.
-Figure 2: Data curation overview.
-(a) The CT image-report pairs are collected from an open-sourced and expert-checked website; (b) and (c) Radiologists provide grounding annotation on any abnormal findings, link to their text description in reports, and categorize into the taxonomy devised by a group of senior radiologists.
-The annotated data are further extended to instruction data with simulated visual prompts and text queries.
 4.2 Annotation Pipeline
@@
 To reduce annotation burden, we adopt a two pre-processing strategy: (i) image-side pre-processing.
-Rather than annotating entire scans, we focus on the key slice and its 8 adjacent slices, as identified and verified by radiologists to be the most informative and representative; (ii) text-side pre-processing.
-We employ GPT-4o gpt4o to extract findings from the reports and identify abnormalities using RaTEScore zhao2024ratescore .
-Here, we recruited 4 clinicians with at least 7 years of radiology experience as annotators.
+Rather than annotating entire scans, we focus on one key slice and its 8 adjacent slices, selected by original uploading radiologists to be the most informative and representative; (ii) text-side pre-processing.
+We employ GPT-4o [gpt4o] to extract findings from the reports and identify abnormalities using RaTEScore [zhao2024ratescore].
+Then, we recruited 4 board-certified radiologists as annotators through a selection phase: candidates each annotated 100 test cases; senior radiologists then selected the top four for full-scale annotation.
 To ensure high-quality annotations, we followed three core principles: (i) representative slices.
@@
 For example, when a report mentions organ enlargement, the finding should not be annotated if enlargement cannot be confidently determined from the available slices.
-Throughout the annotation process, we continuously refined OminiAbnorm-CT-14K based on annotator feedback: (i) taxonomy refinement.
+Throughout the annotation, senior radiologists provided real-time consultation and guidance to resolve uncertain cases from the annotators.
+Meanwhile, we continuously refined OmniAbnorm-CT-14K based on annotator feedback: (i) taxonomy refinement.
 We remove or merge 4 less representative categories, add 22 previously omitted but clinically important ones, and revise 21 categories for improved clarity and expression; (ii) long-tail mitigation.
-To address data imbalance, we identify the underrepresented categories and use GPT-4o to filter scans referencing these rare findings, prioritizing them for annotation.
-As shown in Appendix Section A.6, it has significantly mitigated the long-tail distribution.
-3.4 Extension to Instruction Data
+To address data imbalance, we identify the underrepresented categories and use GPT-4o [gpt4o] to filter scans referencing these rare findings, prioritizing them for annotation.
+More details about the long-tail mitigation can be found in the Appendix.
+4.3 Instruction Data Construction
 As shown in Figure 2(b), we extend our annotated data to support three instruction tasks that simulate realistic clinical scenarios: (i) grounded report generation simulates a comprehensive examination.
@@
 We create visual prompts from annotation masks using various formats, e.g., bounding box, ellipse, contour, and cropped region.
-This simulates a semi-automated workflow where the model elaborates the clinician-identified abnormality in details.
-More details are in Appendix Section A.10.
+This simulates a semi-automated workflow where the model elaborates on the clinician-identified abnormality in detail.
+More details about the instruction data are in the Appendix.
 4.4 Summary
-We prioritize axial CT images for annotation, while also incorporating coronal and sagittal images based on clinical recommendations.
-In total, we annotated 18,969 abnormalities on 9,990 axial scans, 2,738 coronal scans, and 1,803 sagittal scans.
-Representative examples are shown in Figure 1, and the distribution of annotated abnormalities is detailed in Appendix Section A.11.
-Annotation quality is verified by a senior radiologist with 12 years’ experience, with details in Appendix A.5.
-4 OminiAbnorm-CT
-This section presents the details of OminiAbnorm-CT, a novel system for grounded CT image interpretation .
+We prioritize axial images for annotation, while also incorporate coronal and sagittal images based on clinical recommendations.
+In total, we annotated 18,969 abnormalities on 9,990 axial images, 2,738 coronal images, and 1,803 sagittal images.
+Representative examples and the distribution of annotated abnormalities are shown in Figure 1.
+Annotation quality and inter-annotator consistency is verified by a radiologist with 12-year experience, detailed in the Appendix.
+5 OmniAbnorm-CT
+This section presents the details of OmniAbnorm-CT, a novel system for grounded CT image interpretation .
 We start with the problem formulation in Section 5.1, then its architectural details in Section 5.2 and finally the training details in Section 5.3.
+Figure 3: OmniAbnorm-CT.
+We bridge a VLM and a segmentation module, to allow grounding evidence acquisition during the generation of abnormality description, and further enhance its comprehension for flexible usage with text instruction and visual prompts.
 5.1 Problem Formulation
-Given a CT image ($\mathcal{I}\in\mathbb{R}^{H\times W\times D}$), this paper aims to build a model $\Phi_{\theta}(\cdot)$, that enables generating the textual descriptions for abnormal regions, with intermediate groundings when necessary:
+Given a CT image ($\mathcal{I}\in\mathbb{R}^{H\times W\times D}$), we aims to build a model $\Phi_{\theta}(\cdot)$, that generates the textual descriptions for abnormalities, with intermediate groundings when necessary:
 $\{\mathcal{R},\mathcal{S}\}=\Phi_{\theta}(\mathcal{T},\mathcal{I},\mathcal{V}),$ (1)
@@
 Additionally, it also supports visual prompts that allow the model to elaborate on user-highlighted abnormalities, enabling an interactive workflow where users can iteratively refine or correct the grounding results for more precise and context-aware descriptions.
+Model Rubric BLEU-1 BLEU-2 BLEU-3 RaTESc BERTSc METEOR ROUGE-1 ROUGE-L RadG
+Axial (n=2193)
+GPT4o 33.41 15.90 6.22 1.63 39.04 84.57 19.61 18.78 14.55 6.52
+QWen25VL7B(OG) 7.14 5.19 0.98 0.13 14.38 37.14 5.86 5.92 4.47 0.72
+ViP-LLaVA 8.31 15.24 4.50 0.63 33.07 85.12 13.67 17.49 14.69 3.28
+MedDr ✓ 20.82 9.43 2.35 0.33 32.27 78.62 9.01 12.19 10.29 3.67
+LLaVA-Med ✓ 15.41 15.62 4.25 0.55 39.11 84.65 15.72 17.76 13.60 6.13
+BiomedGPT ✓ 20.37 10.85 4.25 0.72 38.29 83.18 15.88 13.50 10.91 4.64
+OmniAbnorm-CT ✓ 35.69 18.03 6.65 1.80 42.81 86.35 19.40 21.66 17.61 9.98
+Coronal (n=750)
+GPT4o 25.41 9.25 3.12 0.92 27.26 59.17 9.25 12.08 9.90 4.26
+QWen25VL7B(OG) 26.25 13.43 3.43 0.65 37.34 84.19 13.34 14.73 11.27 3.34
+ViP-LLaVA 10.27 13.69 3.67 0.51 35.08 84.58 12.54 16.91 13.88 3.68
+MedDr ✓ 26.79 8.22 2.03 0.26 32.98 76.53 8.17 11.71 9.81 4.13
+LLaVA-Med ✓ 22.34 15.60 4.11 0.65 40.96 84.39 14.15 17.99 13.45 6.46
+BiomedGPT ✓ 20.25 12.58 4.79 0.74 38.69 82.90 15.37 14.69 11.46 4.39
+OmniAbnorm-CT ✓ 31.30 18.38 7.19 2.39 42.88 86.00 19.00 21.30 16.87 9.44
+Sagittal (n=591)
+GPT4o 24.39 10.23 3.45 0.88 32.41 72.53 10.01 13.77 11.65 4.00
+QWen25VL7B(OG) 20.09 13.22 3.47 0.59 35.51 83.93 12.82 14.54 11.23 2.95
+ViP-LLaVA 11.56 12.87 3.81 0.38 34.24 84.39 11.86 16.19 13.44 2.84
+MedDr ✓ 19.79 6.89 1.70 0.26 30.32 74.72 6.91 10.07 8.37 2.29
+LLaVA-Med ✓ 15.68 15.01 4.66 0.69 38.97 84.06 14.13 17.64 13.46 5.09
+BiomedGPT ✓ 17.59 12.37 4.73 0.70 36.24 82.54 14.93 14.35 11.22 3.12
+OmniAbnorm-CT ✓ 29.00 17.29 6.48 1.82 41.77 85.62 17.45 20.43 15.87 7.51
+Table 1: Quantitative results on visual prompted report generation.
+Results are averaged within each category and then across all categories, with the best bolded and the second best underlined.
+Models optimized with medical data are marked with ✓.
 5.2 Architecture
-Figure 3: OminiAbnorm-CT.
-We bridge a multi-modal language model and a segmentation module, to allow grounding evidence acquisition during the generation of abnormality description, and further enhance its comprehension for flexible usage with text instruction and visual prompts.
-The overall architecture of OminiAbnorm-CT is illustrated in Figure 3, which comprises a standard multi-modal language model and a segmentation module.
+The overall architecture of OmniAbnorm-CT is illustrated in Figure 3, which comprises a VLM and a segmentation module.
 Our core innovation lies in the integration of them, which enables localizing the abnormality and generating a textual description (findings), supporting flexible interaction through both visual prompts and textual instructions.
-Multi-modal language model functions as the core to comprehend user instruction, invoke the segmentation module properly, and document the abnormal findings based on the grounding evidence.
+Vision language model functions as the core to comprehend user instruction, invoke the segmentation module properly, and document the abnormal findings based on the grounding evidence.
 Specifically, in one workflow to generate grounded reports, given a CT image ($\mathcal{I}$) and text instruction ($\mathcal{T}$) as user input, it will first reason for the specific abnormal findings of interest for the user (or simply any abnormalities).
@@
 $\mathcal{S}=\Phi_{\text{seg}}(\mathcal{I},h_{\text{seg}}),\text{ }h_{\text{seg}}\in\mathbb{R}^{d}$ (2)
-Given the segmentation results ($S$), we convert it to visual prompt ($\mathcal{V}$) in box format, i.e., directly overlay it on the original image to create a new input image ($\mathcal{I}^{\prime}$), and the multi-modal language model could seamlessly continue generating the abnormality description ($\mathcal{R}$) based on it:
+Given the segmentation results ($S$), we convert it to visual prompt ($\mathcal{V}$) in box format, i.e., directly overlay it on the original image to create a new input image ($\mathcal{I}^{\prime}$), and the VLM could seamlessly continue generating the abnormality description ($\mathcal{R}$) based on it:
 $\mathcal{R}=\Phi_{\text{MLLM}}(\mathcal{T},\mathcal{I}^{\prime}),\text{ }\mathcal{I}^{\prime}=\mathcal{I}\oplus\mathcal{V}$ (3)
-where $\oplus$ represents the superimposing operation.
-This leverages the inherent ability of multi-modal language model to perceive visual markers cai2024vip , guiding its attention to specific regions of interest.
+where $\oplus$ represents superimposition. This leverages the inherent ability of VLM to perceive visual markers [cai2024vip], guiding its attention to specific regions of interest.
 Note that, the above design naturally supports another workflow when users have manually delineated specific abnormalities, or intend to refine the grounding results.
-In such case, $\mathcal{V}$ can also be user-input visual prompts on the image, and the multi-modal language model elaborates the user-highlighted abnormality following Equation 3.
+In such case, $\mathcal{V}$ can also be user-input visual prompts on the image, and the VLM elaborates the user-highlighted abnormality following Equation 3.
 Segmentation module is invoked for abnormality grounding based on the CT image ($\mathcal{I}$) and prompt ($h_{\text{seg}}$), summarized as Equation 2.
@@
 where $v$ denotes concatenation of multi-scale image embeddings along the channel dimension, which are down-sampled from each encoder layer to a unified resolution, and $u$ is the pixel-level dense feature from the last decoder layer.
-To bridge the gap between latent spaces, we further align the segmentation prompt ($h_{\text{seg}}$) with the image embeddings ($v$):
+To bridge the gap between latent spaces, we align the segmentation prompt ($h_{\text{seg}}$) with the image embeddings ($v$):
 $q=\Phi_{\text{crossattn}}(v,f(h_{\text{seg}})),\text{ }q\in\mathbb{R}^{d}$ (5)
@@
 5.3 Training
-Training object. We train the multi-modal language model and the segmentation module jointly with a text generation loss and a segmentation loss:
-$\mathcal{L}_{\textbf{txt}}=\text{CE}(\hat{\mathcal{R}},\mathcal{R}),\text{\hskip 6.0pt}\mathcal{L}_{\textbf{seg}}=\text{BCE}(\hat{\mathcal{S}},\mathcal{S})+\text{DICE}(\hat{\mathcal{S}},\mathcal{S})$ (7)
+Training object. We train the VLM and the segmentation module jointly with a text generation loss and a segmentation loss:
+[…]
+$\mathcal{L}_{\textbf{seg}}$ $=\text{BCE}(\hat{\mathcal{S}},\mathcal{S})+\text{DICE}(\hat{\mathcal{S}},\mathcal{S})$ (8)
 Training data.
-Our training data primarily originates from OminiAbnorm-CT-14K, and is formulated into three instruction data formats as illustrated in Section 3.4.
+Our training data primarily originates from OmniAbnorm-CT-14K, and is formulated into three instruction data formats as illustrated in Section 4.3.
 Additionally, we incorporate two types of data as supplements: (i) lesion segmentation data.
 This type of data can be formulated into grounding tasks and utilized to enhance the grounding ability.
-A detailed list of these datasets can be found in Appendix Table 6; (ii) medical VQA data.
+A detailed list of these datasets can be found in the Appendix; (ii) medical visual question-answering (VQA) data.
 This type of data is involved to maintain our model’s generalization capabilities with more diverse question types.
-Specifically, we take CT images and corresponding question-answering data from PubMedVision chen2024huatuogpt .
+Specifically, we take CT images and corresponding question-answering data from PubMedVision [chen2024huatuogpt].
 Implementation details.
-In practice, we adopt QWen2.5-VL-7B yang2024qwen2 as the multi-modal language model, and a 6-layer U-Net as the segmentation module.
+In practice, we adopt QWen2.5-VL-7B [yang2024qwen2] as the VLM, and a 6-layer U-Net as the segmentation module.
 The cross-attention module consists of 6 layers.
-As shown in Appendix Section A.8, we find that expanding the input context range (i.e., using more slices) does not enhance the report generation performance of the multi-modal language model.
-Thus, we only feed the central slice to the model.
-However, for the segmentation module, we include up to 8 adjacent slices as additional input.
-To save computational cost, we only optimize the inserted LoRA layers hu2022lora in the multi-modal language model while leaving the whole segmentation module trainable.
-More implementation details are in Appendix Section A.3.
+We feed 9 consecutive slices to the segmentation module, matching the annotation setup.
+Our collaborating radiologists confirm this window provides sufficient local 3D context for the vast majority of clinically significant findings.
+However, as detailed in the Appendix, expanding the input context (i.e., multiple slices) does not improve VLM report generation.
+We therefore feed only the center slice to the VLM.
+To save computational cost, we only optimize the inserted LoRA layers [hu2022lora] in the VLM while leaving the whole segmentation module trainable.
+More details, runtime and computational analysis are in the Appendix.
 6 Experiments
-In this section, we first present experimental settings in Section 5.1, including task formulation, metrics, baselines, and test datasets.
-We then provide detailed quantitative experimental results in Section 5.2, and ablation study in Section 5.3.
-In addition, the qualitative experimental results are in the Appendix Section A.1 and more ablation studies are in the Appendix A.8 and A.9.
-Table 1: Quantitative results on visual prompted report generation.
-Results are averaged within each category and then across all categories, with the best bolded and the second best underlined.
-Models optimized with medical data are marked with ✓.
-Model B-1 B-2 B-3 RTSc BTSc MTR R-1 R-L RG
+Model DSC Rubric B-1 B-2 B-3 RaTESc BERTSc MTR R-1 R-L RadG
 Axial (n=2193)
-GPT-4o 15.80 6.15 1.59 39.04 84.58 19.31 18.62 14.41 6.51
-QWen2.5-VL-7B 11.57 4.20 0.76 38.35 84.10 16.68 14.26 11.09 3.85
-ViP-LLaVA 15.21 4.49 0.69 33.07 85.15 13.65 17.41 14.68 3.35
-MedFlamingo ✓ 11.26 1.71 0.00 17.35 80.34 9.43 11.69 10.55 0.50
-MedDr ✓ 9.54 2.46 0.41 32.24 78.80 9.13 12.23 10.34 3.74
-LLaVA-Med ✓ 15.55 4.23 0.55 39.08 84.68 15.58 17.67 13.55 6.06
-BiomedGPT ✓ 10.80 4.16 0.71 38.16 83.19 15.77 13.40 10.81 4.65
-OminiAbnorm-CT ✓ 18.03 6.65 1.80 42.81 86.35 19.40 21.66 17.61 9.98
+MedULS + LLaVA-Med 14.92 3.00 10.60 5.31 0.50 26.84 72.92 15.90 12.46 10.50 2.26
+LiSA + LLaVA-Med 20.62 4.87 12.85 6.53 0.70 31.09 81.68 19.11 14.46 12.32 2.75
+BiomedParse + LLaVA-Med 15.69 3.45 13.35 6.86 0.80 31.61 83.04 19.62 14.98 12.77 2.84
+OmniAbnorm-CT ✓ 36.04 13.03 18.50 11.75 7.90 33.24 86.10 21.83 21.65 18.75 3.78
 Coronal (n=750)
-GPT-4o 9.34 3.15 0.96 27.13 58.83 9.27 12.25 10.01 4.31
-QWen2.5-VL-7B 13.62 3.54 0.66 37.26 84.19 13.49 14.97 11.47 3.37
-ViP-LLaVA 13.85 3.81 0.53 35.21 84.60 12.52 17.07 13.96 3.80
-MedFlamingo ✓ 10.98 1.66 0.00 17.73 79.98 9.01 11.60 10.34 0.53
-MedDr ✓ 8.07 2.07 0.29 32.77 76.07 8.09 11.55 9.63 4.08
-LLaVA-Med ✓ 15.65 4.18 0.68 40.99 84.38 14.18 18.15 13.52 6.54
-BiomedGPT ✓ 12.60 4.79 0.70 38.47 82.89 15.45 14.78 11.52 4.45
-OminiAbnorm-CT ✓ 18.38 7.19 2.39 42.88 86.00 19.00 21.30 16.87 9.44
+MedULS + LLaVA-Med 12.73 4.03 9.54 4.91 0.58 28.65 72.77 13.75 12.61 10.68 2.24
+LiSA + LLaVA-Med 17.61 5.66 10.76 5.54 0.68 33.04 81.89 16.04 14.20 11.91 2.94
+BiomedParse + LLaVA-Med 14.77 6.53 11.94 6.07 0.67 32.35 81.29 16.73 15.18 12.64 3.06
+OmniAbnorm-CT ✓ 31.65 13.29 19.13 11.86 7.81 35.47 85.58 20.66 22.14 18.26 4.55
 Sagittal (n=591)
-GPT-4o 10.29 3.40 0.84 32.90 72.76 10.07 13.86 11.63 4.04
-QWen2.5-VL-7B 13.28 3.54 0.63 35.67 83.98 13.01 14.75 11.34 2.91
-ViP-LLaVA 12.84 3.83 0.38 34.51 84.44 11.88 16.29 13.44 2.86
-MedFlamingo ✓ 10.57 1.53 0.00 17.36 78.87 8.66 11.40 10.14 0.36
-MedDr ✓ 6.98 1.77 0.28 30.77 75.35 7.12 10.32 8.54 2.42
-LLaVA-Med ✓ 15.10 4.68 0.66 39.19 84.10 14.18 17.90 13.57 5.11
-BiomedGPT ✓ 12.59 4.81 0.73 36.62 82.60 15.19 14.62 11.35 3.37
-OminiAbnorm-CT ✓ 17.29 6.48 1.82 41.77 85.62 17.45 20.43 15.87 7.51
+MedULS + LLaVA-Med 12.87 2.16 9.71 4.97 0.70 26.06 69.51 13.18 12.26 10.50 1.64
+LiSA + LLaVA-Med 14.82 4.36 12.10 6.16 0.57 31.40 82.77 16.95 14.84 12.76 2.00
+BiomedParse + LLaVA-Med 13.72 4.51 12.25 6.21 0.57 31.90 81.70 16.93 14.98 12.82 2.31
+OmniAbnorm-CT ✓ 34.38 10.11 17.10 10.63 7.10 33.13 85.25 19.05 20.92 17.67 3.70
+Table 2: Quantitative results on grounded report generation. Results are averaged within each category and across all categories, with the best results bolded and the second best underlined.
+In this section, we first illustrate the experimental settings in Section 6.1, including test data, task formulation, and baselines; Then we introduce the evaluation metrics for different tasks in Section 6.2, including existing metrics, and propose AbnormRubric for report generation evaluation; Finally, we provide detailed internal validation results in Section 6.3, and external validation results in Section 6.4.
+Due to limited space, qualitative experiment results, and ablation studies on the segmentation module, context range, and visual prompts are presented in the Appendix.
 6.1 Experiment Settings
-Based on OminiAbnorm-CT-14K, we build a rigorous and comprehensive benchmark to evaluate both grounding and description generation abilities for abnormalities on whole-body CT images.
+Internal validation data.
+We split OmniAbnorm-CT-14K into the train and test set for training and internal validation, following three principles: (i) Axial, coronal, and sagittal images are split independently; (ii) To guarantee a comprehensive benchmark, we allocate at least 5 samples per category to the axial test set (2 for coronal and sagittal), or all samples if fewer exist, with the remaining data following a 3:1 train-test ratio.
+(iii) To avoid data leakage, images from the same patient are restricted to the same set.
+External validation data.
+In addition, we apply the annotation pipeline described in Section 4.2 to 65 chest CT scans sampled from CT-RATE [hamamci2024developing], a large-scale non-contrast chest CT dataset collected in Turkey clinical centers, which serves as an external validation set to assess generalizability.
+Annotators are instructed to label all visible abnormalities in the full volume and, for each abnormality, to select the most prominent slice (i.e., the slice with the largest segmentation area) for evaluation.
+This yields a total of 271 abnormality annotations.
 Task formulation.
@@
 For text-guided grounded report generation, we generate queries for both present and absent abnormalities at a 3:1 ratio to simulate realistic diagnostic scenarios.
-Metrics.
-For grounding results, we calculate Dice Similarity Coefficient (DSC); for generated descriptions, we calculate BLEU (B) bleu , RaTEScore (RTSc) zhao2024ratescore , BERTScore (BTSc) zhangbertscore , METEOR (MTR) banerjee2005meteor , ROUGE (R) rouge and RadGraph (RG) delbrouck2024radgraph .
-In the text-guided grounded report generation, we prompt models to output a specific response when the queried abnormalities are confirmed absent.
-Specificity (SP) is then calculated as the proportion of correctly identified absent abnormalities out of all truly absent abnormalities in the test set.
-Meanwhile, when models miss the queried abnormalities, both DSC and all description metrics are assigned as 0.
 Baseline.
-For visual prompted report generation, we compare with multi-modal LLMs in both the medical domain (MedDr meddr , BiomedGPT zhang2024generalist , LLaVA-Med li2023llava , and Med-Flamingo moor2023med ) and the general domain (GPT-4o gpt4o , Qwen2.5-VL yang2024qwen2 , and ViP-LLaVA cai2024vip ).
+For visual prompted report generation, we compare with multi-modal LLMs in both the medical domain (MedDr [meddr], BiomedGPT [zhang2024generalist], and LLaVA-Med [li2023llava]) and the general domain (GPT-4o [gpt4o], Qwen2.5-VL [yang2024qwen2], and ViP-LLaVA [cai2024vip]).
 For (text-guided) grounded report generation, as no such solution exists in the literature, we prompt LLaVA-Med to describe abnormal findings based on segmentation results from auxiliary grounding models.
-Since the images in OminiAbnorm-CT-14K lack the original DICOM metadata, we only consider 2D segmentation models as baselines, including BiomedParse zhao2024foundation , LiSA lai2024lisa , and MedULS de2024uls23 .
-More details are in Appendix Section A.4.
-Data.
-We independently split the train and test data for axial, coronal, and sagittal images.
-To guarantee diversity, we allocate at least 5 samples per category to the axial test set (2 for coronal and sagittal), or all samples if fewer exist, with the remaining data following a 3:1 train-test ratio.
-To avoid data leakage, images from the same patient are restricted to the same set.
-5.2 Experiment Results
-Table 2: Quantitative results on grounded report generation. Results are averaged within each category and across all categories, with the best results bolded and the second best underlined.
-Model DSC B-1 B-2 B-3 RTSc BTSc MTR R-1 R-L RG
+Since the images in OmniAbnorm-CT-14K lack the original DICOM metadata, we only consider 2D segmentation models as baselines, including BiomedParse [zhao2024foundation], LiSA [lai2024lisa], and MedULS [de2024uls23].
+More details are in the Appendix.
+Model DSC Rubric B-1 B-2 B-3 RaTESc BERTSc MTR R-1 R-L RadG
 Axial (n=2193)
-LiSA + LLaVA-Med 20.96 12.94 6.65 0.73 31.38 82.07 19.20 14.62 12.44 2.79
-BiomedParse + LLaVA-Med 15.78 13.61 6.98 0.84 31.50 82.36 19.84 15.28 13.02 2.74
-MedULS + LLaVA-Med 15.24 12.70 6.43 0.74 31.05 84.27 18.74 14.67 12.42 2.31
-OminiAbnorm-CT 36.04 18.50 11.75 7.90 33.24 86.10 21.83 21.65 18.75 3.78
+MedULS + LLaVA-Med 14.24 3.58 9.27 2.32 0.27 26.58 78.54 11.06 11.03 8.86 1.78
+LiSA + LLaVA-Med 17.34 0.25 1.78 0.41 0.25 4.40 22.11 2.03 2.34 2.08 0.24
+BiomedParse + LLaVA-Med 16.42 0.56 1.91 0.44 0.19 4.73 23.34 2.15 2.54 2.22 0.27
+OmniAbnorm-CT ✓ 32.40 14.82 11.94 3.33 0.74 34.80 83.85 12.94 14.65 11.94 4.32
 Coronal (n=750)
-LiSA + LLaVA-Med 17.09 11.48 5.85 0.64 33.28 82.27 16.23 14.36 12.12 2.74
-BiomedParse + LLaVA-Med 14.55 11.71 5.95 0.68 32.89 81.80 16.88 14.75 12.57 3.35
-MedULS + LLaVA-Med 12.77 11.64 5.88 0.64 32.25 84.11 16.45 14.71 12.34 2.96
-OminiAbnorm-CT 31.65 19.13 11.86 7.81 35.47 85.58 20.66 22.14 18.26 4.55
+MedULS + LLaVA-Med 12.21 3.30 10.08 2.42 0.17 27.81 78.85 10.72 11.57 9.06 1.34
+LiSA + LLaVA-Med 12.45 0.10 2.18 0.48 0.24 5.05 28.04 2.47 2.99 2.67 0.27
+BiomedParse + LLaVA-Med 14.41 0.29 2.25 0.45 0.22 5.46 27.92 2.49 3.02 2.65 0.27
+OmniAbnorm-CT ✓ 27.74 20.29 12.51 4.10 1.22 36.56 82.98 13.61 15.83 12.36 5.72
 Sagittal (n=591)
-LiSA + LLaVA-Med 14.72 11.10 5.60 0.58 31.51 81.12 16.18 14.07 11.96 2.18
-BiomedParse + LLaVA-Med 14.45 10.89 5.46 0.52 32.10 79.37 15.68 14.06 11.95 2.13
-MedULS + LLaVA-Med 12.95 10.95 5.32 0.50 30.94 83.53 15.39 14.12 11.83 1.87
-OminiAbnorm-CT 34.38 17.10 10.63 7.10 33.13 85.25 19.05 20.92 17.67 3.70
+MedULS + LLaVA-Med 11.61 3.35 10.07 2.60 0.32 26.87 78.33 10.61 11.83 9.16 1.79
+LiSA + LLaVA-Med 11.36 0.36 2.06 0.43 0.29 5.05 24.51 2.23 2.66 2.15 0.34
+BiomedParse + LLaVA-Med 14.25 0.27 2.28 0.66 0.31 5.95 29.05 2.71 3.28 2.81 0.36
+OmniAbnorm-CT ✓ 29.89 18.37 12.80 4.51 1.42 36.04 83.97 14.02 16.20 12.79 4.85
+Table 3: Quantitative results on text-guided grounded report generation. Results are averaged within each category and then across all categories, with the best bolded and the second best underlined.
+6.2 Evaluation Metrics
+Existing Metrics.
+For grounding results, we calculate Dice Similarity Coefficient (DSC).
+For generated report, we include natural language generation metrics: BLEU (B) [bleu], BERTScore (BERTSc) [zhangbertscore], METEOR (MTR) [banerjee2005meteor], and ROUGE (R) [rouge]; as well as medical-specific metrics: RaTEScore (RaTESc) [zhao2024ratescore] and RadGraph (RadG) [delbrouck2024radgraph].
+AbnormRubric (Rubric) is proposed as a metric for generated report.
+Unlike above metrics that are limited to simple word matching or unexplainable semantic similarity, AbnormRubric aligns with clinical judgment by comprehensively evaluating the abnormalities in the generated report from: (i) the detection ratio, (ii) the description accuracy of radiological attributions, and (iii) the hallucination.
+Formally, given all the ground-truth abnormalities ($\text{TP}+\text{FN}$), a LLM is first instructed to numerate the detected (TP) and hallucinated abnormalities (FP) in the generated report:
+$\text{Recall}=\text{TP}/(\text{TP}+\text{FN})$ (9)
+$\text{Precision}=\text{TP}/(\text{TP}+\text{FP})$ (10)
+Then, for each detected abnormality, the LLM is instructed to judge whether the generated description align with ground-truth report on different radiological attributions:
+$\text{Accuracy}=\frac{1}{N}\sum_{i=1}^{N}\mathbb{I}[\text{Attribution}_{i}\text{ aligns with GT}]$ (11)
+where $\mathbb{I}$ is the indicator function.
+Notably, for each type of abnormality in our defined taxonomy, the radiologists have tailored and identified a set of radiological attributions, including location, size, morphology, and density, that are clinically important and should be accurately described.
+And the final score is formulated as:
+$\text{Recall}^{*}=(1+\text{Accuracy})/2\times\text{Recall}$ (12)
+$\text{Rubric}=2\times\frac{\text{Recall}^{*}\times\text{Precision}}{\text{Recall}^{*}+\text{Precision}}$ (13)
+More details of AbnormRubric are in the Appendix.
+6.3 Internal Validation Results
 Visual Prompted Report Generation.
-As depicted in Table 1, OminiAbnorm-CT consistently outperforms all baselines across all planes and metrics.
-Specifically, compared to the strongest baseline LLaVA-Med, OminiAbnorm-CT achieves average improvements of 2.5 in BLEU-1, 2.7 in RaTEScore, and 3.1 in RadGraph across all three planes.
-These results demonstrate OminiAbnorm-CT’s effectiveness in semi-automated workflows, where radiologists identify abnormalities and the model provides accurate and detailed interpretations.
+Table 1 reports that OmniAbnorm-CT outperforms all baselines on 29 out of 30 metrics.
+Specifically, compared to the strongest baseline, OmniAbnorm-CT achieves average improvements of 2.4 in BLEU-1, 2.8 in RaTEScore and 4.26 in AbnormRubric across all three planes.
+It demonstrates OmniAbnorm-CT’s effectiveness in semi-automated workflows, where radiologists mark the abnormalities and the model provides accurate and detailed interpretations.
 Grounded Report Generation.
-As shown in Table 2, OminiAbnorm-CT demonstrates superior grounding capabilities, outperforming the best baselines by 15.08, 14.56, and 19.66 in DSC across axial, coronal, and sagittal planes respectively.
-For report generation, OminiAbnorm-CT achieves the best performance across all metrics, with notable improvements in both RadGraph and RaTEScore across all anatomical planes, indicating that the generated reports can better align with clinical standards.
-These results demonstrate OminiAbnorm-CT’s potential for fully automated abnormality localization and report generation in clinical settings.
-Table 3: Quantitative results on text-guided grounded report generation. Results are averaged within each category and then across all categories, with the best bolded and the second best underlined.
-Model DSC SP B-1 B-2 B-3 RTSc BTSc MTR R-1 R-L RG
-Axial (n=2193)
-MedULS + LLaVA-Med 14.53 6.10 9.54 2.18 0.19 27.35 82.98 11.45 11.48 9.20 1.61
-LiSA + LLaVA-Med 17.48 21.35 10.57 2.41 0.28 29.18 83.47 11.59 12.34 9.80 1.61
-BiomedParse + LLaVA-Med 16.57 19.87 10.66 2.38 0.26 29.66 83.39 11.74 12.62 9.86 2.16
-OminiAbnorm-CT 32.40 58.43 11.94 3.33 0.74 34.80 83.85 12.94 14.65 11.94 4.32
-Coronal (n=750)
-MedULS + LLaVA-Med 12.26 5.37 10.73 2.60 0.26 29.34 82.89 11.40 12.79 9.82 1.61
-LiSA + LLaVA-Med 12.04 13.76 11.58 2.71 0.22 30.76 83.21 11.36 13.40 10.23 1.83
-BiomedParse + LLaVA-Med 13.60 17.79 11.21 2.64 0.25 30.24 83.33 11.47 13.50 10.31 2.34
-OminiAbnorm-CT 27.74 45.16 12.51 4.10 1.22 36.56 82.98 13.61 15.83 12.36 5.72
-Sagittal (n=591)
-MedULS + LLaVA-Med 11.89 7.66 10.25 2.69 0.29 27.40 82.49 11.24 12.17 9.53 1.23
-LiSA + LLaVA-Med 11.23 16.67 11.14 2.72 0.30 29.79 82.88 11.15 13.36 10.07 1.42
-BiomedParse + LLaVA-Med 15.35 17.57 11.07 2.72 0.39 31.12 83.01 11.66 13.85 10.40 1.78
-OminiAbnorm-CT 29.89 45.41 12.80 4.51 1.42 36.04 83.97 14.02 16.20 12.79 4.85
+As shown in Table 2, OmniAbnorm-CT demonstrates superior grounding capabilities, outperforming the best baselines by 15.42, 14.04, and 19.56 in DSC on axial, coronal, and sagittal planes respectively.
+For report generation, OmniAbnorm-CT achieves the best performance across all metrics, with notable improvements in AbnormRubric, RadGraph and RaTEScore across all anatomical planes, indicating that the generated reports can better align with clinical standards.
+These results demonstrate OmniAbnorm-CT’s potential for fully automated abnormality localization and report generation in clinical settings.
 Text-guided Grounded Report Generation.
-As shown in Table 3, OminiAbnorm-CT demonstrates superior grounding capabilities, with DSC scores improving by 14.24, 14.14, and 14.54 in axial, coronal, and sagittal planes respectively compared to the best baselines.
-For report generation, OminiAbnorm-CT excels across both lexical overlap metrics (average improvements of 1.3 on BLEU-1 and 2.2 on ROUGE-1 across all planes) and clinical relevance metrics (average gains of 4.0 on RaTEScore and 2.9 on RadGraph across all planes), confirming its superior ability to generate reports that align with clinical assessments.
-Additionally, OminiAbnorm-CT achieves significantly higher Specificity, demonstrating its robustness to queries about non-existing abnormalities.
-5.3 Ablation Study
-The integration of the segmentation module allows OminiAbnorm-CT to acquire grounding evidence for abnormal findings and generate descriptions accordingly.
-To validate its necessity, we conduct an ablation study comparing the full model against a variant where the segmentation module is removed (denoted as OminiAbnorm-CT w/o Seg).
-This modified version directly generates findings descriptions without anatomical localization guidance.
-In the text-guided grounded report generation task, as shown in Table 4, OminiAbnorm-CT without the segmentation module demonstrates notable performance degradation, with inferior results on 23 of 27 metrics across three planes.
-Furthermore, detailed analysis reveals that without the segmentation module, the model struggles to accurately detect the queried abnormalities in CT images, exhibiting a pronounced negative bias that leads to high specificity scores.
-Similarly, in the grounded report generation task, as depicted in Table 5, removing the segmentation module leads to consistent performance degradation on all the metrics, further confirming that segmentation-based evidence is crucial for generating high-quality medical reports.
-Table 4: Ablation study of the segmentation module for text-guided grounded report generation task. Results are averaged within each category and then across all categories, with the best bolded
-Model SP B-1 B-2 B-3 RTSc BTSc MTR R-1 R-L RG
-Axial (n=2193)
-OminiAbnorm-CT w/o Seg 95.39 4.50 2.33 1.21 20.53 31.06 7.04 8.50 7.02 7.25
-OminiAbnorm-CT 58.43 11.94 3.33 0.74 34.80 83.85 12.94 14.65 11.94 4.32
-Coronal (n=750)
-OminiAbnorm-CT w/o Seg 95.71 4.50 2.47 1.27 19.07 28.34 6.27 7.96 6.43 6.17
-OminiAbnorm-CT 45.16 12.51 4.10 1.22 36.56 82.98 13.61 15.83 12.36 5.72
-Sagittal (n=591)
-OminiAbnorm-CT w/o Seg 92.08 4.17 2.16 1.12 16.62 25.43 5.84 7.06 5.53 5.67
-OminiAbnorm-CT 45.41 12.80 4.51 1.42 36.04 83.97 14.02 16.20 12.79 4.85
-Table 5: Ablation study of the segmentation module in grounded report generation task. Results are averaged within each category and then across all categories, with the best bolded
-Model B-1 B-2 B-3 RTSc BTSc MTR R-1 R-L RG
-Axial (n=2193)
-OminiAbnorm-CT w/o Seg 7.73 4.74 3.16 27.87 73.14 17.65 15.18 13.58 3.42
-OminiAbnorm-CT 19.00 12.26 8.40 33.85 86.24 22.45 22.27 19.43 4.60
-Coronal (n=750)
-OminiAbnorm-CT w/o Seg 8.12 4.81 3.12 31.03 75.55 15.64 15.27 13.23 3.84
-OminiAbnorm-CT 19.28 12.03 7.96 35.40 85.56 20.81 22.35 18.47 4.65
-Sagittal (n=591)
-OminiAbnorm-CT w/o Seg 8.79 5.21 3.33 31.21 77.41 16.18 15.86 13.87 3.91
-OminiAbnorm-CT 16.69 10.41 6.98 33.45 84.94 18.77 20.67 17.61 3.83
-6 Related Work
-Medical image segmentation aims to delineate clinically meaningful regions on medical images.
-For radiology images, most datasets focus on annotations for important organs and anatomical structures Totalsegmentator ; abdomenatlas ; FLARE ; AMOS22 or lesions heller2021kidney ; MSD ; BraTS2023GLI ; de2024uls23 ; yan2018deep .
-Specialist segmentation models nnUNET ; UNet ; zhou2023nnformer ; zhou2019unet++ have achieved remarkable success in the past decade, while recently, building foundational generalist models MedSAM ; zhao2024foundation ; zhao2023one ; du2024segvol has received increased attention.
-However, few works contribute to the grounding of abnormal image findings, which encompass a broader spectrum including abnormal organs, potential lesions, and more generally, any anomalies meaningful for clinical decision-making.
-Medical image report generation aims to faithfully interpret the medical images with formal text, which is indispensable for numerous clinic procedures rao2025multimodal .
-Existing datasets with image-report pairs are limited to specific modalities and regions, such as chest X-ray johnson2019mimic , brain MRI lei2024autorg and chest CT hamamci2024foundation ; chen2024bimcv , overlooking the clinical demands for multi-plane whole-body CT image report generation.
-Recent progress zhang2024radgenome ; chen2025large ; lei2024autorg ; chen2024vision on grounded report generation are limited to structured organ-level segmentation and description, failing to localize and interprete fine-grained abnormalities.
-Generative visual-language foundation models for medicine are built on large-scale multi-modal medical data, and have demonstrated exceptional performance and generalization capabilities across diverse medical tasks wu2023towards ; moor2023med ; zhang2024generalist ; li2023llava ; meddr .
-Despite their potential application to multi-plane, whole-body CT images, however, most existing approaches fail to simultaneously provide grounded evidence during generation, leading to poor explainability and hallucinations that are more difficult to detect.
+In Table 3, OmniAbnorm-CT achieves substantially better grounding results, improving DSC over the strongest baseline on axial (+14.24), coronal (+14.14), and sagittal (+14.54).
+For report generation, OmniAbnorm-CT outperforms in both general overlap metrics (average +1.3 BLEU-1, +2.2 ROUGE-1 across planes) and clinical metrics (average +14.42 AbnormRubric across planes), indicating closer alignment with clinical assessments.
+It indicate that OmniAbnorm-CT can effectively follows radiologist instructions to ground and interpret the relevant abnormalities.
+Model DSC Rubric B-1 B-2 B-3 RaTESc BERTSc MTR R-1 R-L RadG
+Visual prompted report generation (n=271)
+GPT4o - 19.15 17.37 7.59 1.81 37.86 85.52 20.46 19.16 15.18 3.73
+QWen25VL7B(OG) - 15.15 6.41 2.57 0.42 37.04 82.84 14.62 9.52 7.37 2.51
+ViP-LLaVA - 10.31 17.59 6.20 0.55 32.88 85.46 15.33 19.18 16.36 1.95
+MedDr ✓ - 13.52 3.24 1.13 0.17 20.36 54.28 3.58 4.94 4.29 0.74
+LLaVA-Med ✓ - 12.22 17.63 7.40 1.20 34.96 85.52 19.81 19.01 15.64 2.92
+BiomedGPT ✓ - 12.07 10.99 4.57 0.60 35.14 82.62 17.30 13.92 11.37 2.67
+OmniAbnorm-CT ✓ - 24.09 19.09 7.71 1.67 39.87 86.64 19.46 21.02 17.34 4.27
+Grounded report generation (n=271)
+MedULS + LLaVA-Med 11.63 3.96 12.86 7.09 1.32 33.53 81.88 17.31 15.88 12.87 2.50
+LiSA + LLaVA-Med 18.34 4.37 12.33 6.54 0.62 35.43 83.57 17.03 15.87 12.71 2.69
+BiomedParse + LLaVA-Med 11.78 5.32 11.98 6.51 0.74 34.76 83.47 16.59 16.10 13.08 2.87
+OmniAbnorm-CT ✓ 22.20 8.62 18.55 12.10 8.14 34.93 86.09 21.61 22.36 18.32 2.57
+Text-guided grounded report generation (n=271)
+MedULS + LLaVA-Med 11.69 3.23 9.47 2.86 0.33 25.57 74.55 11.57 10.81 8.91 1.32
+LiSA + LLaVA-Med 9.72 0.17 2.85 1.43 0.89 4.68 20.26 3.08 3.32 2.90 1.40
+BiomedParse + LLaVA-Med 10.50 0.33 4.33 2.06 1.44 6.65 25.52 4.50 4.77 4.48 1.63
+OmniAbnorm-CT ✓ 14.13 15.12 12.38 3.76 1.18 32.25 76.34 13.55 14.85 11.97 3.55
+Table 4: External validation results across three tasks.
+Metrics are averaged within each abnormality category and then across all categories.
+Within each block, the best result is bolded and the second best is underlined.
+For metrics that are not applicable to a given task, we report “-”.
+Models optimized with medical data are marked with ✓.
+6.4 External Validation Results
+As demonstrated in Table 4, on the external validation set, OmniAbnorm-CT maintains strong performance across all three tasks.
+In visual prompted report generation, OmniAbnorm-CT achieves the highest AbnormRubric score (+4.94 over the best baseline), indicating that its generated reports align more closely with clinical criteria.
+Moreover, in both grounded report generation and text-guided grounded report generation, OmniAbnorm-CT surpasses all baselines on report generation metrics and shows superior grounding ability (DSC improvements of +10.57 for grounded report generation and +2.44 for text-guided grounded report generation).
+These results suggest that OmniAbnorm-CT generalizes well to external datasets.
 7 Conclusion
 This paper advances the automatic grounded interpretation of CT imaging from an abnormality-centric view.
-To support this, we develop a hierarchical taxonomy of 404 representative abnormal findings on CT images.
-We contribute OminiAbnorm-CT-14K, a meticulously annotated dataset with detailed grounding and description annotation for around 19K abnormalities, from 14.5K multi-plane CT images across the entire human body.
-Built on this dataset, OminiAbnorm-CT enables automatic grounding and description of abnormalities, while supporting flexible clinical usage with text queries or visual prompts.
-Comprehensive evaluations across three clinically representative tasks demonstrate OminiAbnorm-CTs superior performance.
-Our work pioneers abnormality-centric CT interpretation, enhancing diagnostic transparency through accurate grounding and detailed characterization of findings, with significant potential to transform radiological practice.
-Acknowledgments and Disclosure of Funding
-Weidi would like to acknowledge the funding from Scientific Research Innovation Capability Support Project for Young Faculty (ZY-GXQNJSKYCXNLZCXM-I22).
-Appendix A Appendix
-A.1 Qualitative Experiment Results
+To support this, we develop a hierarchical taxonomy of 404 representative abnormal findings.
+We contribute OmniAbnorm-CT-14K, a meticulously annotated dataset with detailed grounding and description annotation for around 19K abnormalities, from 14.5K multi-plane CT images across the entire human body.
+Built on this dataset, OmniAbnorm-CT enables automatic grounding and description of abnormalities, while supporting flexible clinical usage with text queries or visual prompts.
+We establish three representative tasks and devise a clinically grounded metric to demonstrate OmniAbnorm-CTs superior performance.
+Limitations.
+As an early attempt, this paper certainly has limitations and future works to continue: On dataset, we annotate representative slices rather than full volumes, primarily due to the high cost of expert labeling.
+We will explore semi-automated tools to scale to full-volume annotations in future.
+For OmniAbnorm-CT, current VLM has limited volumetric perception, constrained by training data and compute resources.
+We plan to curate larger-scale whole-body CT datasets for pretraining and adapting a 3D vision encoder in future.
+Nevertheless, our work pioneers abnormality-centric CT interpretation, improving diagnostic transparency via accurate grounding and detailed characterization of abnormalities.
+We believe this paradigm shift holds substantial potential to transform radiology practice.
+Rethinking Whole-Body CT Image Interpretation:
+An Abnormality-Centric Approach
+Appendix
+Appendix A Qualitative Experiment Results
 Fig.
-4 presents several examples from the grounded report generation task, comparing the segmentation and report generation results between OminiAbnorm-CT and BiomedParse+LLaVA-Med.
-The results clearly demonstrate that BiomedParse fails to detect any abnormalities in the latter two cases, consequently leading LLaVA-Med to generate reports that are hardly relevant to the images.
-In contrast, OminiAbnorm-CT successfully localizes the abnormal regions across all examples and produces comparatively more accurate reports for the abnormalities.
+4 presents several cases from the grounded report generation task, comparing the segmentation and report generation results between OmniAbnorm-CT and BiomedParse+LLaVA-Med.
+It clearly shows that BiomedParse fails to detect any abnormalities in the latter two cases, and consequently, LLaVA-Med generates reports irrelevant to the abnormalities.
+In contrast, OmniAbnorm-CT successfully localizes the abnormalities across all cases and produces more accurate reports.
 Figure 4: Qualitative comparison on the grounded report generation task.
 In Fig.
-5, we demonstrate several examples from the text-guided grounded report generation task, and compare the segmentation and report generation results of OminiAbnorm-CT and BiomedParse+LLaVA-Med.
-As illustrated in the first two examples, BiomedParse fails to accurately localize the queried abnormalities.
-While in the latter two examples, LLaVA-Med is unable to correctly interpret the segmentation results when generating reports.
-In contrast, OminiAbnorm-CT can successfully localize the queried abnormalities across all examples and generates more precise and clinically relevant reports.
+5, we compare OmniAbnorm-CT with BiomedParse+LLaVA-Med on the text-guided grounded report generation task.
+BiomedParse mislocalizes the queried findings in first two cases, while LLaVA-Med fails to correctly interpret the segmentation results in last two cases.
+In contrast, OmniAbnorm-CT consistently localizes the queried abnormalities and produces more precise, clinically aligned reports.
 Figure 5: Qualitative comparison on the text-guided grounded report generation task.
 In Fig.
-6, we demonstrate several representative examples in the visual prompted report generation task, comparing the abnormality descriptions generated by OminiAbnorm-CT, LLaVA-Med, and QWen2.5-VL-7B.
-It shows that LLaVA-Med and QWen2.5-VL-7B exhibit some fundamental errors in their results.
-For instance, in the first case, LLaVA-Med misidentifies the kidney as the gallbladder; While in the last case, QWen2.5-VL-7B fails to distinguish between the liver and gallbladder.
-By contrast, OminiAbnorm-CT is able to precisely capture and identify the abnormal findings highlighted by the visual prompts and generate high quality reports.
+6, we compare OmniAbnorm-CT with LLaVA-Med and QWen2.5-VL-7B on visual prompted report generation.
+The baselines show some fundamental mistakes, e.g., LLaVA-Med mislabels the kidney as the gallbladder in the first case, and QWen2.5-VL-7B confuses the liver and gallbladder in the last case.
+In contrast, OmniAbnorm-CT accurately identifies the marked abnormalities and produces higher-quality, clinically consistent reports.
 Figure 6: Qualitative comparison on the visual prompted report generation task.
-A.2 Limitations
-As a pioneering attempt toward automatic grounding and interpretation of abnormalities on multi-plane and whole-body CT images, our work certainly has several limitations: (i) Regarding data annotation, despite providing a large-scale dataset, some rare abnormalities (55 out of 404 in the taxonomy) remain uncovered, as shown in Section A.11.
-Additionally, many abnormalities have very few annotations, which significantly restricts the model’s capability range.
-Expanding annotations for these rare categories is treated as our future work.
-Meanwhile, the CT images in OminiAbnorm-CT-14K may also exhibit demographic and regional bias.
-(ii) Due to resource constraints, we currently provide grounding annotation for a few representative slices on each CT image, rather than full volumetric segmentation.
-In the future, we could potentially address this limitation by using interactive segmentation models such as nnInteractive [27].
-(iii) Limited by the 3D image perception capability of multi-modal language models, which is also a common challenge in current generative visual-language foundation models for medicine [53, 44, 58, 33, 20], we only considered single slice as input when generating reports.
-Enhancing models’ ability to reason across adjacent slices and integrate volumetric context represents a critical direction for our future research.
-A.3 More Implementation Details of OminiAbnorm-CT
+Appendix B Ablation Studies
+We conduct a series of ablation studies on the key factors of OmniAbnorm-CT, from the segmentation module (Section B.1), the context range of input image (Section B.2), to the form of visual prompts (Section B.3).
+This section introduces the settings and results of these experiments sequentially.
+B.1 Ablation Study on Segmentation Module
+The integration of the segmentation module allows OmniAbnorm-CT to acquire grounding evidence for abnormal findings and generate descriptions accordingly.
+To validate its necessity, we conduct an ablation study comparing the full model against a variant without the segmentation module (OmniAbnorm-CT w/o Seg), which directly generates findings.
+Text-guided grounded report generation.
+As shown in Table 5, removing the segmentation module significantly degrades performance on most metrics (23/27 across planes).
+Furthermore, our manual analysis reveals that, without the grounding evidence, the model struggles to accurately detect the queried abnormalities in CT images.
+Grounded report generation.
+Similarly, Table 6 shows that OmniAbnorm-CT w/o Seg consistently underperforms the full model on all metrics, reaffirming that segmentation-driven evidence is essential for accurate grounding and generating clinically faithful reports.
+Model BLEU-1 BLEU-2 BLEU-3 RaTESc BERTSc METEOR ROUGE-1 ROUGE-L RadG
+Axial (n=2193)
+OmniAbnorm-CT w/o Seg 4.50 2.33 1.21 20.53 31.06 7.04 8.50 7.02 7.25
+OmniAbnorm-CT 11.94 3.33 0.74 34.80 83.85 12.94 14.65 11.94 4.32
+Coronal (n=750)
+OmniAbnorm-CT w/o Seg 4.50 2.47 1.27 19.07 28.34 6.27 7.96 6.43 6.17
+OmniAbnorm-CT 12.51 4.10 1.22 36.56 82.98 13.61 15.83 12.36 5.72
+Sagittal (n=591)
+OmniAbnorm-CT w/o Seg 4.17 2.16 1.12 16.62 25.43 5.84 7.06 5.53 5.67
+OmniAbnorm-CT 12.80 4.51 1.42 36.04 83.97 14.02 16.20 12.79 4.85
+Table 5: Ablation study of the segmentation module for text-guided grounded report generation task.
+Model BLEU-1 BLEU-2 BLEU-3 RaTESc BERTSc METEOR ROUGE-1 ROUGE-L RadG
+Axial (n=2193)
+OmniAbnorm-CT w/o Seg 7.73 4.74 3.16 27.87 73.14 17.65 15.18 13.58 3.42
+OmniAbnorm-CT 19.00 12.26 8.40 33.85 86.24 22.45 22.27 19.43 4.60
+Coronal (n=750)
+OmniAbnorm-CT w/o Seg 8.12 4.81 3.12 31.03 75.55 15.64 15.27 13.23 3.84
+OmniAbnorm-CT 19.28 12.03 7.96 35.40 85.56 20.81 22.35 18.47 4.65
+Sagittal (n=591)
+OmniAbnorm-CT w/o Seg 8.79 5.21 3.33 31.21 77.41 16.18 15.86 13.87 3.91
+OmniAbnorm-CT 16.69 10.41 6.98 33.45 84.94 18.77 20.67 17.61 3.83
+Table 6: Ablation study of the segmentation module in grounded report generation task.
+B.2 Ablation Study on Context Range
+In clinical practice, adjacent-slice context is important for CT interpretation, including segmentation and diagnosis [nnUNET, zheng2024large].
+To assess its impact on report generation, we vary the number of input slices in visual prompted report generation, using a base QWen2.5-VL-7B.
+Surprisingly, Table 7 shows that adding slices yields no consistent gains across metrics.
+We hypothesize that current LVLMs, which encode the slices independently into tokens, struggle to model subtle inter-slice spatial relations.
+Accordingly, we use only the center slice as input to the LVLM in subsequent training and evaluation of OmniAbnorm-CT.
+Input Slices BLEU-1 BLEU-2 BLEU-3 RaTESc BERTSc METEOR ROUGE-1 ROUGE-L RadGraph
+Axial (n=2193)
+4 Adjacent Slices 9.39 2.98 0.45 31.12 83.52 14.90 11.51 9.01 2.46
+2 Adjacent Slices 9.24 3.03 0.46 31.47 83.49 15.05 11.41 8.99 2.49
+No Adjacent Slices 11.57 4.20 0.76 38.35 84.10 16.68 14.26 11.09 3.85
+Coronal (n=750)
+4 Adjacent Slices 10.93 3.93 0.85 32.95 83.49 15.82 13.29 9.79 3.62
+2 Adjacent Slices 10.89 3.85 0.73 33.06 83.48 15.87 13.34 9.85 3.71
+No Adjacent Slices 13.62 3.54 0.66 37.26 84.19 13.49 14.97 11.47 3.37
+Sagittal (n=591)
+4 Adjacent Slices 9.39 2.98 0.45 31.12 83.52 14.90 11.51 9.01 2.46
+2 Adjacent Slices 11.48 4.16 0.94 33.86 83.32 15.70 13.64 9.97 3.75
+No Adjacent Slices 13.28 3.54 0.63 35.67 83.98 13.01 14.75 11.34 2.91
+Table 7: Impact of multi-slice context range on visual prompted report generation, with the base Qwen2.5-VL-7B model as baseline.
+B.3 Ablation Study on Different Visual Prompts
+Recent work shows VLMs respond exhibit varying perceptual capabilities for different prompt formats [cai2024vip].
+We therefore study visual prompt design for visual prompted report generation.
+For each abnormality, we simulate four prompts, i.e., center crop, ellipse, contour, and box, and evaluate OmniAbnorm-CT separately.
+We also consider a per-case oracle that takes the best score among the four, representing the theoretical best performance via selecting the optimal prompt for each specific abnormality, depending on its shape, location, and etc.
+As shown in Table 8, cropping performs worst on most metrics (24/27), likely due to lost context.
+Ellipse, contour, and box yield similar performance.
+The oracle notably improves all metrics over any single prompt, revealing large per-case variance.
+This suggests that adaptively choosing the visual prompt by morphology and location is beneficial, whereas cropping away context is harmful.
+Prompt Type BLEU-1 BLEU-2 BLEU-3 RaTEScore BERTScore METEOR ROUGE-1 ROUGE-L RadGraph
+Axial (n=2193)
+Center Cropping 11.70 2.87 0.68 33.03 84.84 12.77 14.58 11.75 3.92
+Ellipse 12.43 3.60 1.15 35.14 84.87 13.67 15.55 12.63 5.27
+Contour 12.60 3.79 1.22 35.74 85.09 14.12 15.97 12.92 5.81
+Bounding Box 12.35 3.69 1.26 35.46 84.92 13.65 15.76 12.80 5.63
+Max 18.03 6.65 1.80 42.81 86.35 19.40 21.66 17.61 9.98
+Coronal (n=750)
+Center Cropping 10.94 2.96 0.80 33.57 84.48 11.92 14.23 11.36 3.87
+Ellipse 12.49 3.79 1.13 36.09 84.56 13.44 15.94 12.43 5.16
+Contour 12.61 3.73 1.02 35.50 84.71 13.45 15.75 12.25 4.87
+Bounding Box 12.15 3.70 1.12 35.12 84.38 13.27 15.68 12.13 4.68
+Max 18.38 7.19 2.39 42.88 86.00 19.00 21.30 16.87 9.44
+Sagittal (n=591)
+Center Cropping 10.98 2.73 0.39 32.23 84.20 11.98 14.45 11.01 3.05
+Ellipse 11.37 3.10 0.65 34.36 84.01 12.19 14.45 10.95 3.46
+Contour 11.08 2.95 0.57 34.19 84.22 12.01 14.38 11.10 3.22
+Bounding Box 11.29 2.98 0.67 34.07 84.13 12.16 14.84 11.24 3.45
+Max 17.29 6.48 1.82 41.77 85.62 17.45 20.43 15.87 7.51
+Table 8: Comparison of different visual prompts in visual prompted report generation task, with OmniAbnorm-CT fixed as baseline.
+Appendix C Details of AbnormRubric
+As introduced in Section 6.2, the computation of AbnormRubric is decomposed into three stages that explicitly mirror the way radiologists assess a report: (i) abnormality detection (recall), (ii) hallucination checking (precision), and (iii) description accuracy of radiological attributions.
+All three stages are implemented using LLM-based judgments given the ground-truth report and the generated report.
+Detection of ground-truth abnormalities (Recall).
+Given the set of abnormalities in the ground-truth report, we iterate over each and query an LLM to determine whether this abnormality is explicitly present in the generated report.
+Abnormalities that are judged as present are counted as true positives (TP), while the remaining ground-truth abnormalities are counted as false negatives (FN).
+The detailed prompt is:
+You are a medical imaging report analyst.
+Your task is to evaluate an AI-generated radiology report (or only a portion/fragment of a report) from two aspects: 1.
+Whether the report is a VALID radiology report (it makes a clear statement about what abnormalities are present OR absent on the current image).
+2.
+Whether the report identifies a specific abnormality.
+ABNORMALITY TO CHECK: Target Abnormality: {abnormality_category} Relevant Description in Detailed: {groundtruth_description} AI-GENERATED REPORT TO ANALYZE: {generated_report} EVALUATION CRITERIA: 1.
+Validity Check: The generated report may contain medically irrelevant content, nonsensical phrases or garbled text.
+Completely ignore such content and focus only on medically relevant content.
+The generated report may be blank, or entirely composed of non-medical gibberish, irrelevant text, or insufficient to determine presence/absence of abnormalities.
+Then ignore the following procedure and your evaluation is complete: is_invalid_report=1, detected=0, description=None.
+2.
+Detection: Report as detected=1 if the target abnormality ({abnormality_category}) is reported in the generated report.
+Consider synonyms and clinically equivalent terms as matches.
+Regardless of the description accuracy, such as incorrect location, size measurements, morphological characteristics, density/attenuation, and so on.
+The generated report may indicate that no abnormalities were detected on the image, for example, I don’t see any relevant abnormalities on the image.
+Then ignore the following procedure and your evaluation is complete: detected=0.
+3.
+Description Extraction: If detected=1, collect ALL relevant description from the report.
+Preserve original wording without modifications.
+If multiple descriptions exist, concatenate them.
+May include adjacent text if the description cannot be grammatically isolated.
+REQUIRED OUTPUT FORMAT:
+  "is_invalid_report": 0, // 1 or 0
+  "detected": 1,      // 1 if detected, 0 if not
+  "description": "exact text" // string if detected, null if not
+STRICT PROHIBITIONS: NEVER add information not present in the generated report. NEVER paraphrase or summarize original text. DO NOT output any additional text except for a valid json.
+Identification of hallucinated abnormalities (Precision).
+To assess hallucinations, we further instruct the LLM to enumerate any abnormalities that are described in the generated report but do not exist in the ground-truth report.
+Abnormalities that cannot be matched are counted as false positives (FP).
+The detailed prompt is:
+Your task is to evaluate the hallucinations in an AI-generated radiology report (or only a portion/fragment of a report): 1.
+Suppose some ground truth abnormalities have been successfully detected by the AI-generated radiology report.
+2.
+Identify any ADDITIONAL abnormalities mentioned in the generated report that are NOT covered by the ground truth list.
+THE AI-GENERATED REPORT TO ANALYZE: generated_report GROUND TRUTH ABNORMALITIES AND MATCHING RESULTS: groundtruth_abnormalities INSTRUCTIONS: 1.
+Review matched content: The above shows which ground truth abnormalities were detected and their corresponding text in the generated report.
+2.
+Find additional abnormalities: Identify ALL ADDITIONAL abnormal findings mentioned in the generated report that are NOT already matched to ground truth abnormalities.
+3.
+Exclude already matched content: Do NOT include any abnormalities that are already matched to ground truth abnormalities.
+4.
+Ignore normal findings: Only report abnormalities, not normal findings.
+5.
+Output: ONLY valid JSON with no additional text.
+NOTE: 1.
+The generated report may contain medically irrelevant content, nonsensical phrases or garbled text.
+Completely ignore such content and focus only on medically relevant content.
+2.
+If the generated report is entirely composed of non-medical gibberish, irrelevant text, or is blank, then your evaluation is complete: count = 0 and hallucinated_abnormalities = [].
+REQUIRED OUTPUT FORMAT:
+  "hallucinated_abnormalities": [...],  // list of hallucinated abnormalities
+  "count": x  // number of hallucinated abnormalities
+If no additional abnormalities found:
+  "hallucinated_abnormalities": [],
+  "count": 0
+Accuracy of radiological attributions.
+For each detected ground-truth abnormality (TP), we assess how accurately its key radiological attributions are described in the generated report.
+Based on our abnormality taxonomy, radiologists define a set of clinically important attributions for each abnormality type (e.g., location, morphology, density, size).
+For a given abnormality, we consider only those attributions that are explicitly mentioned in the ground-truth report and ask the LLM to compare them with the corresponding text in the generated report.
+For each attribution, the LLM returns a binary judgment of whether the generated description is clinically consistent with the ground-truth description.
+The detailed prompt is:
+You are a medical imaging report analyst.
+For a single abnormality in the image, the AI model has generated a description.
+Your task is to compare the AI-generated description with the ground-truth description and evaluate the AI’s accuracy across specific attributes that will be provided.
+EVALUATION INSTRUCTIONS: INPUT DATA: Abnormality: {abnormality_category} Ground-Truth (GT) Description: {gt_description} AI-Generated (Pred) Description: {pred_description} Attributes: {attribution_ls}
+1.
+Detection of Attribute Mention: Evaluate ONLY the pre-defined attributes provided in the list.
+For each attribute, check if it is mentioned in the GT description.
+For attributes present in the GT description, check if they are also mentioned in the Pred description.
+2.
+Comparison: For attributes present in BOTH GT and Pred: Extract the exact phrasing from both descriptions.
+Determine if the descriptions are clinically equivalent (1) or not (0).
+Consider synonyms and clinically equivalent terms as matches.
+Provide brief reasoning for your judgment.
+For attributes ONLY present in GT: Mark as not clinically equivalent (0).
+3.
+Quantitative Data Handling: If GT contains quantitative data (e.g., measurements with units): Extract the numerical value and unit from GT.
+Check if Pred provides quantitative data for the same attribute.
+If Pred lacks quantitative data: mark as non-equivalent (0).
+If Pred provides data: calculate percentage error after unit conversion.
+Error $\leq 20\%$: equivalent (1); Error $20\%$: non-equivalent (0).
+In the reasoning field, include the brief calculation process.
+4.
+Special Cases: If Pred provides quantitative data but GT does not: mark as non-equivalent (0).
+For ambiguous cases where clinical equivalence cannot be determined: mark as uncertain (1-1).
+5.
+Strict Constraints: Base ALL judgments strictly on the provided GT and Pred texts.
+NO additions, deletions, modifications, or external inferences.
+Preserve exact wording when extracting descriptions.
+Clinical relevance overrides literal text matching: Ignore differences in writing style, terminology preferences, and phrasing variations.
+Consider clinically synonymous terms as equivalent.
+Focus on clinical semantic meaning rather than exact word matching.
+Account for standard medical abbreviations and their equivalent expressions.
+REQUIRED OUTPUT FORMAT (JSON only):
+  "attributes": {
+    "attribute_name_1": {
+      "present_in_gt": 1, // 0 if not present, and the followings can be empty
+      "gt_description": "exact text from GT",  // null if not present
+      "pred_description": "exact text from Pred",  // null if not present
+      "equivalent": 1,  // 0 if not equivalent, -1 if can not be determined
+      "reasoning": "brief clinical justification",  // null if not present
+      "gt_quantified": 1,  // 0 if not quantified
+      "pred_quantified": 1,  // 0 if not quantified
+      "error_percentage": 5.2 // null if not both quantified
+      or can not be calculated
+Appendix D Supplementary Implementation Details
+D.1 Supplementary Implementation Details of OmniAbnorm-CT
 Training Hyperparameters.
 We adopt 8 NVIDIA A100-80G GPUs for training, with 1 sample per device and gradient accumulation set to 4.
-We train OminiAbnorm-CT for 250K iterations in total, and use 5% steps for warming up.
-We take AdamW [36] as optimizer with a learning rate of 2e-4.
+We train OmniAbnorm-CT for 250K iterations in total, and use 5% steps for warming up.
+We take AdamW [loshchilov2017decoupled] as optimizer with a learning rate of 2e-4.
 When optimizing the multi-modal language model and the segmentation module jointly, we assign equal weights to the text generation loss and segmentation loss, as defined in Equation 8.
@@
 Training Data.
-We pad and rescale all images to 512$\times$512.
+We pad and rescale all images to $\times512$.
 For training data sampling from four tasks: visual prompted report generation, grounded report generation, text-guided grounded report generation, and general VQA, we control the ratio to be 1:1:1:1.
-Additionally, in the abnormality grounding task, we maintain a ratio of 6:4 between OminiAbnorm-CT-14K and public lesion segmentation data.
+Additionally, in the abnormality grounding task, we maintain a ratio of 6:4 between OmniAbnorm-CT-14K and public lesion segmentation data.
 For each abnormality description, we use the following prompt to have GPT rewrite three different versions, then randomly select one as the ground truth:
@@
 Input: Original medical report written by a doctor Output: THREE rewritten versions with identical content but different expression Please rewrite the following medical report in three different ways: \$Findings Description\$ Your response must follow this exact format: \$Output Template\$
-A.4 Implementation Details of Baselines
-In visual prompted generation task, we prompt each abnormality with all types of visual prompts, and take the maximum score. All the methods are prompted with the following template:
+D.2 Supplementary Implementation Details of Baselines
+Visual prompted generation. we prompt each abnormality with all types of visual prompts, and take the maximum score. All the methods are prompted with the following template:
 You are a helpful medical assistant.
@@
 Do not list general possibilities, explanations, or recommendations.
-In grounded report generation task, we integrate LLaVA-Med with 2D segmentation models as baselines.
-Specifically, we re-implement the MedULS [13] with a 2D nnU-Net [26] and the public lesion segmentation datasets in Table 6, covering all the datasets in the official ULS-23 challenge and 10 additional ones.
+Grounded report generation.
+We integrate LLaVA-Med with 2D segmentation models as baselines.
+Specifically, we re-implement the MedULS [de2024uls23] with a 2D nnU-Net [nnUNET] and the public lesion segmentation datasets in Table 9, covering all the datasets in the official ULS-23 challenge and 10 additional ones.
 To our knowledge, this represents the segmentation model with the broadest capability range (in terms of abnormality variety) that can be constructed from currently available public datasets.
-For BiomedParse [61] and LiSA [31], we prompt them with ‘Abnormal findings on the CT image’ to derive segmentation results for all the abnormalities on the input CT image.
+For BiomedParse [zhao2024foundation] and LiSA [lai2024lisa], we prompt them with ‘Abnormal findings on the CT image’ to derive segmentation results for all the abnormalities on the input CT image.
 Then, we combine each segmentation model with LLaVA-Med by converting the segmentation results into bounding boxes overlaid on the CT image, and prompt LLaVA-Med to generate the report based on these visual cues:
@@
 Respond with ’I don’t see any abnormalities on the image.’ if no abnormalities are present.
-In text-guided grounded report generation task, we use the simulated text queries as prompts for BiomedParse and LiSA to derive segmentation results for the queried abnormalities, which are detailed in Section 3.4.
+Text-guided grounded report generation.
+We use the simulated text queries as prompts for BiomedParse and LiSA to derive segmentation results for the queried abnormalities, which are detailed in Section 4.3.
 For MedULS, since it doesn’t support text-prompted segmentation, we simply use its unconditioned segmentation results.
@@
 Respond with ’I don’t see any relevant abnormalities on the image.’ if no abnormalities are present.
-A.5 Quality Verification on OminiAbnorm-CT-14K
-We conducted rigorous and comprehensive quality verification on the annotations in OminiAbnorm-CT-14K.
+D.3 Runtime and Computation Analysis
+Training.
+We use 4 A100-80G GPUs for training OmniAbnorm-CT.
+The segmentation module is first pre-trained for 50,000 steps with batch size 16, which takes around 10 hours.
+Then the Qwen2.5-VL is fine-tuned on the visual-prompted generation task for 20,000 steps, with batch size of 1 and gradient accumulation every 8 steps, which takes around 8 hours.
+Finally, the VLM and segmentation modules are jointly trained for grounded report generation, with batch size of 1 and gradient accumulation every 8 steps.
+This final stage takes approximately 48 hours for 100,000 steps.
+Inference.
+All inferences are conducted on 1 A100-80G GPU.
+We use mixed-precision without quantization.
+The average inference latency is 0.49 second per sample for visual-prompted generation, 0.98 second per sample for grounded report generation, and 1.01 second per sample for text-guided grounded report generation.
+Appendix E Supplementary Details of OmniAbnorm-CT-14K
+E.1 Quality Verification
+Annotation Quality.
+We conducted rigorous and comprehensive quality verification on the annotations in OmniAbnorm-CT-14K.
 For each annotator, we randomly sampled 100 annotated images on the axial plane, 50 on the coronal, and 50 on the sagittal plane.
 Then a senior radiologist with 12 years of experience assesses the annotation quality on four key metrics: (i) Detection rate measures the percentage of abnormalities successfully identified and annotated by the annotator, without any omission; (ii) Grounding precision evaluates the percentage of grounding annotations that properly encompass the primary regions of the abnormality, while minimizing false positive areas; (iii) Report concordance quantifies the percentage of description annotations that faithfully reflect the linked abnormal findings on the image; (iv) Classification accuracy measures the percentage of abnormalities that are correctly categorized.
-Our verification results demonstrated that: (i) 3 out of 4 annotators achieved perfect scores across all four metrics (100%).
+The verification results demonstrated that: (i) 3 out of 4 annotators achieved perfect scores across all four metrics (100%).
 (ii) Only one annotator had 10 abnormality annotations where the descriptions were not sufficiently accurate and the category labels were also incorrect, resulting in Report concordance and Classification accuracy of 95%, while maintaining 100% Detection rate and Grounding precision.
-These results confirm the high quality of annotations in our OminiAbnorm-CT-14K.
-A.6 The Long-Tail Distribution of OminiAbnorm-CT-14K
-To mitigate the long-tail distribution in OminiAbnorm-CT-14K, we identified underrepresented organs in our annotated corpus and strategically employed GPT-4o to analyze unlabeled reports.
+These results confirm the high quality of annotations in our OmniAbnorm-CT-14K.
+Cross-annotator Consistency.
+We further conducted inter-annotator consistency checks across 4 annotators on 40 random samples.
+For each sample, we enumerated all annotator pairs and, for each pair, computed the DSC between their segmentation masks and the BLEU-1 score between their abnormality-level report descriptions.
+Averaging these metrics across all annotator pairs yielded mean pairwise scores of DSC=84.5 for segmentation and BLEU-1=72.2 for finding descriptions, indicating reasonably good agreement among the annotators.
+E.2 Mitigation of the Long-Tail Distribution
+To mitigate the long-tail distribution in OmniAbnorm-CT-14K, we identified underrepresented organs in our annotated corpus and strategically employed GPT-4o to analyze unlabeled reports.
 Using the following prompt, we efficiently filtered cases containing abnormal findings related to these underrepresented organs:
@@
 (a) Comparison of organ distribution in annotations; (b) Comparison of abnormality category distribution in annotations.
-A.7 Comparison between OminiAbnorm-CT-14K and Existing Datasets
-We compare OminiAbnorm-CT-14K with existing public CT image datasets in Table 6, including those widely used for lesion segmentation, lesion detection, organ segmentation, or report generation.
-In contrast to the substantial limitations exhibited by these datasets, as detailed in Sec.
-3.1, OminiAbnorm-CT-14Kepresents the first large-scale dataset designed for abnormality grounding and description across multi-plane and whole-body CT images.
-Table 6: Comparison of key characteristics between OminiAbnorm-CT-14K and widely-used public CT imaging datasets.
+E.3 Comparison against Existing Datasets
+We compare OmniAbnorm-CT-14K with existing public CT image datasets in Table 9, including those widely used for lesion segmentation, lesion detection, organ segmentation, or report generation.
+In contrast to the substantial limitations exhibited by these datasets, as detailed the main paper, OmniAbnorm-CT-14Kepresents the first large-scale dataset designed for abnormality grounding and description across multi-plane and whole-body CT images.
+Dataset Task Anatomy Plane Report #Category #Image
+ULS Bone [de2024uls23] Lesion Seg. Bone Axial
+1 151
+ULS Pancreas [de2024uls23] Lesion Seg. Pancreas Axial
+1 119
+MSD Liver [MSD] Lesion Seg. Liver Axial
+1 131
+MSD Lung [MSD] Lesion Seg. Lung Axial
+1 63
+MSD Colon [MSD] Lesion Seg. Colon Axial
+1 126
+MSD Pancreas [MSD] Lesion Seg. Pancreas Axial
+1 281
+COVID19 [COVID19] Lesion Seg. Lung Axial
+1 20
+KiTS23 [KiTS23] Lesion Seg. Kidney Axial
+2 489
+KiPA22 [KiPA22] Lesion Seg. Kidney Axial
+2 70
+NSCLC [NSCLC] Lesion Seg. Lung Axial
+1 85
+LIDC IDRI [LIDCIDRI] Lesion Seg. Lung Axial
+1 750
+LNDb [LNDb] Lesion Seg. Lung Axial
+1 236
+INSTANCE22 [INSTANCE] Lesion Seg. Brain Axial
+1 100
+Seg.Rap2023 [SegRap2023] Lesion Seg. Head & Neck Axial
+2 120
+FUMPE [FUMPE] Lesion Seg. Lung Axial
+1 35
+RibFrac [ribfracchallenge2025] Lesion Seg. Rib Axial
+1 420
+Adrenal ACC Ki67 [AdrenalACCKi67] Lesion Seg. Adrenal Gland Axial
+1 29
+LNQ2023 [dorent2024lnq] Lesion Seg. Lung Axial
+1 393
+NIH-LN [roth2014new] Lesion Seg. Lung Axial
+1 175
+CCC-18 [CCC18] Lesion Seg. Chest & Abdomen Axial
+404
+DeepLesion [yan2018deep] Lesion Det. Whole Body Axial
+33K
+TotalSegmentator [Totalsegmentator] Organ Seg. Whole Body Axial
+1.2K
+AbdomenAtlas [abdomenatlas] Organ Seg. Chest & Abdomen Axial
+8K
+CTRATE [haaga2016ct] Report Gen. Chest Axial
+26K
+BIMCV-R [chen2024bimcv] Report Gen. Chest Axial
+8K
+ReXGroundingCT [baharoon2025rexgroundingct] Lesion Seg. Chest Axial 14 3K
+Lesion Seg. Axial 340 10K
+OmniAbnorm-CT-14K Lesion Det. Whole Body Coronal 255 2K
+Report Gen. Sagittal 223 1.6K
+Table 9: Comparison of key characteristics between OmniAbnorm-CT-14K and widely-used public CT imaging datasets.
 Note that CCC18 and DeepLesion has no category label for each annotated lesion.
 Even though some scans in these datasets are isotropic, they are acquired as axial-plane.
-Abbreviations: Rpt = Report; Cat = Abnormality Category; Img = Image; Seg.
+Abbreviations: Category = Abnormality Category; Seg.
 = Segmentation; Det.
 = Detection; Gen = Generation.
-Dataset Task Anatomy Plane Rpt #Cat #Img
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-Lesion Seg. Axial 340 10K
-OminiAbnorm-CT-14K Lesion Det. Whole Body Coronal 255 2K
-Report Gen. Sagittal 223 1.6K
-A.8 Ablation Study on Context Range
-In clinical practice, the contextual information provided by adjacent slices is critical for the interpretation of CT images, which is also widely validated in tasks such as medical image segmentation [26] and diagnosis [64].
-To investigate the impact of context range on report generation, in the visual prompted report generation task, we adjust the number of input slices and evaluate its impact on a base QWen2.5-VL-7B model.
-As shown in Table 7, when we increase the input slices, there is no consistent performance improvement on any metric.
-We speculate this is mainly because multi-modal language models, which generally encode and project each input image into tokens independently, can hardly capture the subtle spatial relationships across adjacent slices.
-Therefore, we only take the center slice as image input to the multi-modal language model in the subsequent training and evaluation of OminiAbnorm-CT.
-Table 7: Impact of multi-slice context range on visual prompted report generation, with the base Qwen2.5-VL-7B model as baseline.
-Abbrevations: B = BLEU, RTSc = RaTEScore, BTSc = BERTScore, MTR = METEOR, R = Rouge, RG = RadGraph.
-Input Slices B-1 B-2 B-3 RTSc BTSc MTR R-1 R-L RG
-Axial (n=2193)
-4 Adjacent Slices 9.39 2.98 0.45 31.12 83.52 14.90 11.51 9.01 2.46
-2 Adjacent Slices 9.24 3.03 0.46 31.47 83.49 15.05 11.41 8.99 2.49
-No Adjacent Slices 11.57 4.20 0.76 38.35 84.10 16.68 14.26 11.09 3.85
-Coronal (n=750)
-4 Adjacent Slices 10.93 3.93 0.85 32.95 83.49 15.82 13.29 9.79 3.62
-2 Adjacent Slices 10.89 3.85 0.73 33.06 83.48 15.87 13.34 9.85 3.71
-No Adjacent Slices 13.62 3.54 0.66 37.26 84.19 13.49 14.97 11.47 3.37
-Sagittal (n=591)
-4 Adjacent Slices 9.39 2.98 0.45 31.12 83.52 14.90 11.51 9.01 2.46
-2 Adjacent Slices 11.48 4.16 0.94 33.86 83.32 15.70 13.64 9.97 3.75
-No Adjacent Slices 13.28 3.54 0.63 35.67 83.98 13.01 14.75 11.34 2.91
-A.9 Ablation Study on Different Visual Prompts
-Recent studies have revealed that VLMs exhibit varying perceptual capabilities for different forms of visual prompts [8].
-Therefore, in the visual-prompted report generation task, we investigated the impact of different visual prompt formats to give insight on optimal design choices in application.
-Specifically, for each abnormality in our dataset, we simulated four distinct visual prompt types: center cropping, ellipse, contour, and bounding box, as input and evaluated OminiAbnorm-CT’s generated reports respectively.
-We also take the maximum score among them for each abnormality, representing the theoretical best performance achievable by selecting the optimal visual prompt type for each specific abnormality, depending on its shape, location, and etc.
-As shown in Table 8, cropping the prompted area consistently underperforms across nearly all metrics (24 out of 27 total measurements), likely due to the loss of crucial contextual information outside the prompted region.
-The other individual visual prompting methods (ellipse, contour, and bounding box) demonstrate comparable performance.
-Notably, integrating multiple visual prompting techniques yields a substantial improvement across all metrics compared to the best individual prompt for each measure.
-This indicates significant performance variation when using different visual prompts for the same abnormality, suggesting that selecting appropriate prompt types based on abnormality morphology and location could be an effective strategy.
-Table 8: Comparison of different visual prompt formats in visual prompted report generation task, with OminiAbnorm-CT fixed as baseline.
-Abbrevations: B = BLEU, RTSc = RaTEScore, BTSc = BERTScore, MTR = METEOR, R = Rouge, RG = RadGraph.
-Prompt Type B-1 B-2 B-3 RTSc BTSc MTR R-1 R-L RG
-Axial (n=2193)
-Center Cropping 11.70 2.87 0.68 33.03 84.84 12.77 14.58 11.75 3.92
-Ellipse 12.43 3.60 1.15 35.14 84.87 13.67 15.55 12.63 5.27
-Contour 12.60 3.79 1.22 35.74 85.09 14.12 15.97 12.92 5.81
-Bounding Box 12.35 3.69 1.26 35.46 84.92 13.65 15.76 12.80 5.63
-Max 18.03 6.65 1.80 42.81 86.35 19.40 21.66 17.61 9.98
-Coronal (n=750)
-Center Cropping 10.94 2.96 0.80 33.57 84.48 11.92 14.23 11.36 3.87
-Ellipse 12.49 3.79 1.13 36.09 84.56 13.44 15.94 12.43 5.16
-Contour 12.61 3.73 1.02 35.50 84.71 13.45 15.75 12.25 4.87
-Bounding Box 12.15 3.70 1.12 35.12 84.38 13.27 15.68 12.13 4.68
-Max 18.38 7.19 2.39 42.88 86.00 19.00 21.30 16.87 9.44
-Sagittal (n=591)
-Center Cropping 10.98 2.73 0.39 32.23 84.20 11.98 14.45 11.01 3.05
-Ellipse 11.37 3.10 0.65 34.36 84.01 12.19 14.45 10.95 3.46
-Contour 11.08 2.95 0.57 34.19 84.22 12.01 14.38 11.10 3.22
-Bounding Box 11.29 2.98 0.67 34.07 84.13 12.16 14.84 11.24 3.45
-Max 17.29 6.48 1.82 41.77 85.62 17.45 20.43 15.87 7.51
-A.10 Generation of Text Queries and Visual Prompts
+E.4 Generation of Text Queries and Visual Prompts
 We simulate four visual prompts for each annotated abnormality, mimicking how clinicians would select the most appropriate highlighting method based on the abnormality’s shape and location: (i) cropped region.
@@
 The information about all the abnormalities on the CT images: \$Abnormal Findings\$ The clinical presentation of the patient corresponding to this CT image is: \$Presentation\$ The overall report for the CT image containing these abnormalities is: \$Whole Report\$ The doctors’ discussion results for the patient corresponding to this CT image are: \$Impression\$ Your response must follow this exact format: \$Output Template\$
-A.11 Detailed Distribution of Abnormalities in OminiAbnorm-CT-14K
-Figure 8: Distribution of OminiAbnorm-CT-14K across 82 anatomical structures and 40 major systems or organs.
-Darker blue indicates higher sample density.
-Abd.
-& Pelv.
-= Abdominal and Pelvic Cavity; Test., Epid., Scrot.
-= Testes, Epididymis, Scrotum; Morph.
-= Morphology; Pos.
-= Positional; Vent.
-& Subarach.
-= Ventricles and Subarachnoid.
-Table 9: Detailed distribution of abnormality categories in OminiAbnorm-CT-14K.
+E.5 Detailed Distribution of OmniAbnorm-CT-14K
 Organ Anatomical Structure Category Axial Coronal Sagittal
@@
 Others 1 0 0
+Table 10: Detailed distribution of abnormality categories in OmniAbnorm-CT-14K.
+Organ Anatomical Structure Category Axial Coronal Sagittal
 Spine Spinal Cord Spinal cord compression 3 0 2
@@
 Others 2 1 5
-Table 10: (Continued) Detailed distribution of abnormality categories in OminiAbnorm-CT-14K.
-Organ Anatomical Structure Category Axial Coronal Sagittal
 Eye Eyeball Eyeball atrophy 4 1 2
@@
 Others 1 1 0
-Middle Ear Middle ear fluid or hemorrhage 5 0 0
+Table 11: (Continued) Detailed distribution of abnormality categories in OmniAbnorm-CT-14K.
+Organ Anatomical Structure Category Axial Coronal Sagittal
+Ear Middle Ear Middle ear fluid or hemorrhage 5 0 0
 Middle ear soft tissue mass 23 23 3
@@
 Others 6 2 0
-Table 11: (Continued) Detailed distribution of abnormality categories in OminiAbnorm-CT-14K.
-Organ Anatomical Structure Category Axial Coronal Sagittal
 Pharynx Pharynx (including nasopharynx, oropharynx, and hypopharynx) Pharyngeal narrowing or obstruction 4 2 1
@@
 Pharyngeal space Pharyngeal space soft tissue mass 46 23 16
+Pharyngeal space cystic mass 13 6 2
+Pharyngeal emphysema 7 0 0
+Others 5 7 1
+Table 12: (Continued) Detailed distribution of abnormality categories in OmniAbnorm-CT-14K.
+Organ Anatomical Structure Category Axial Coronal Sagittal
+Pharynx Pharyngeal space Pharyngeal space soft tissue mass 46 23 16
 Pharyngeal space cystic mass 13 6 2
@@
 Others 3 0 1
-Table 12: (Continued) Detailed distribution of abnormality categories in OminiAbnorm-CT-14K.
+Table 13: (Continued) Detailed distribution of abnormality categories in OmniAbnorm-CT-14K.
 Organ Anatomical Structure Category Axial Coronal Sagittal
@@
 Others 17 1 0
-Myocardium Myocardial hypertrophy 4 1 1
+Table 14: (Continued) Detailed distribution of abnormality categories in OmniAbnorm-CT-14K.
+Organ Anatomical Structure Category Axial Coronal Sagittal
+Heart Myocardium Myocardial hypertrophy 4 1 1
 Myocardial thinning 3 0 0
@@
 Others 4 0 0
-Table 13: (Continued) Detailed distribution of abnormality categories in OminiAbnorm-CT-14K.
-Organ Anatomical Structure Category Axial Coronal Sagittal
-Heart Coronary arteries Coronary artery myocardial bridge 1 0 0
+Coronary arteries Coronary artery myocardial bridge 1 0 0
 Coronary artery dilation 0 0 0
@@
 Others 0 0 0
+Table 15: (Continued) Detailed distribution of abnormality categories in OmniAbnorm-CT-14K.
+Organ Anatomical Structure Category Axial Coronal Sagittal
 Gastrointestinal tract (including esophagus, stomach, and intestines) Gastrointestinal lumen Gastrointestinal dilatation 265 212 112
@@
 Others 0 4 2
-Table 14: (Continued) Detailed distribution of abnormality categories in OminiAbnorm-CT-14K.
-Organ Anatomical Structure Category Axial Coronal Sagittal
 Liver Hepatic parenchyma Hepatic parenchymal morphological alteration 99 34 11
@@
 Others 3 1 0
-Gallbladder wall Gallbladder wall thickening 53 11 1
+Table 16: (Continued) Detailed distribution of abnormality categories in OmniAbnorm-CT-14K.
+Organ Anatomical Structure Category Axial Coronal Sagittal
+Gallbladder Gallbladder wall Gallbladder wall thickening 53 11 1
 Gallbladder wall calcification 9 1 0
@@
 Others 17 1 0
-Table 15: (Continued) Detailed distribution of abnormality categories in OminiAbnorm-CT-14K.
+Table 17: (Continued) Detailed distribution of abnormality categories in OmniAbnorm-CT-14K.
 Organ Anatomical Structure Category Axial Coronal Sagittal
@@
 Others 4 3 0
+Table 18: (Continued) Detailed distribution of abnormality categories in OmniAbnorm-CT-14K.
+Organ Anatomical Structure Category Axial Coronal Sagittal
+Bladder Bladder cavity Bladder distention 4 0 1
+Bladder stone 24 5 3
+Bladder content density change 13 2 1
+Others 4 3 0
 Bladder wall Bladder wall diffuse thickening 17 5 6
@@
 Others 1 0 2
-Table 16: (Continued) Detailed distribution of abnormality categories in OminiAbnorm-CT-14K.
-Organ Anatomical Structure Category Axial Coronal Sagittal
 Adrenal gland Adrenal gland Adrenal soft tissue mass 122 17 1
@@
 Others 0 0 0
+Testes, epididymis, and scrotum Testis Testicular enlargement 1 0 0
+Testicular atrophy 0 0 0
+Testicular soft tissue mass 1 0 0
+Testicular calcification 0 0 0
+Testicular cystic mass 0 0 0
+Testicular torsion 0 0 0
+Testicular hemorrhage and rupture 0 0 0
+Others 0 0 0
+Table 19: (Continued) Detailed distribution of abnormality categories in OmniAbnorm-CT-14K.
+Organ Anatomical Structure Category Axial Coronal Sagittal
 Testes, epididymis, and scrotum Testis Testicular enlargement 1 0 0
@@
 Others 4 1 1
-Table 17: (Continued) Detailed distribution of abnormality categories in OminiAbnorm-CT-14K.
-Organ Anatomical Structure Category Axial Coronal Sagittal
 Fallopian tube Fallopian tube Fallopian tube thickening 3 0 0
@@
 Others 0 0 0
+Table 20: (Continued) Detailed distribution of abnormality categories in OmniAbnorm-CT-14K.
+Organ Anatomical Structure Category Axial Coronal Sagittal
 Ovary Ovary Ovarian enlargement 9 3 0
@@
 Others 30 50 26
-Table 18: (Continued) Detailed distribution of abnormality categories in OminiAbnorm-CT-14K.
+Table 21: (Continued) Detailed distribution of abnormality categories in OmniAbnorm-CT-14K.
 Organ Anatomical Structure Category Axial Coronal Sagittal
@@
 Others 17 3 8
+Table 22: (Continued) Detailed distribution of abnormality categories in OmniAbnorm-CT-14K.